### v1

# Data-driven Analysis of First-Order Methods via Distributionally Robust Optimization

Abstract
We consider the problem of analyzing the probabilistic performance of first-order methods when solving convex optimization problems drawn from an unknown distribution only accessible through samples. By combining performance estimation (PEP) and Wasserstein distributionally robust optimization (DRO), we formulate the analysis as a tractable semidefinite program. Our approach unifies worst-case and average-case analyses by incorporating data-driven information from the observed convergence of first-order methods on a limited number of problem instances. This yields probabilistic, data-driven performance guarantees in terms of the expectation or conditional value-at-risk of the selected performance metric. Experiments on smooth convex minimization, logistics regression, and Lasso show that our method significantly reduces the conservatism of classical worst-case bounds and narrows the gap between theoretical and empirical performance.
1 Introduction
First-order methods have gained significant attention over the past decade due to their low computational cost per iteration, modest memory requirements, and ability to warm-start from previous solutions [4, 60]. These advantages make them particularly valuable for large-scale problems in data science and machine learning, as well as embedded optimization applications in engineering and optimal control.
These algorithms have been around since the 1950s [23], and have been extended to handle nondifferentiable functions and constraints via proximal operators [54], and composite functions via operator splitting techniques such as the alternating directions method of multipliers (ADMM) [10], or primal-dual hybrid gradient method (PDHG) [12]. Recently, several general-purpose solvers based on first-order methods appeared in the literature, including PDLP [1, 2] for linear programs (LPs), OSQP [67] for quadratic programs (QPs), and SCS [52, 51] and COSMO [25] for semidefinite programs (SDPs). Despite these advances, the behavior of these algorithms are still highly sensitive to data, often exhibiting slow convergence for badly-scaled problems.
To design faster and more reliable first-order methods, a central question is to understand the accuracy of the solutions returned. In most convergence analyses, this translates into estimating the worst-case accuracy after a certain number of iterations for a given function class. While most of these results focus on asymptotic convergence rates [6, 49], a recent line of work called the performance estimation problem (PEP) [21, 69], focuses on constructing finite-step guarantees by solving a semidefinite program (SDP). By defining the appropriate constraints, i.e., interpolation conditions, this technique allows us to represent various popular convex function classes (e.g., strongly convex, smooth, Lipschitz, indicators) and virtually all first-order methods used in practice. This approach has led to several new results in the optimization community including tighter convergence rate proofs [68, 73], and even faster optimization algorithms [35, 55, 32].
However, the practical convergence behavior of these schemes can be very different from the best known classical or computer-assisted worst-case bounds. In some cases, first-order methods might converge significantly faster, as outlined in the following example.
Example 1 (Motivating example).
Consider a least-squares problem with box constraints:
$\begin{array}[]{ll}\textnormal{minimize}&f(x)=(1/2)\|Ax-b\|^{2}\\ \textnormal{subject to}&0\leq x\leq 1,\end{array}$ (1)
where $A\in{\mathbf{R}}^{n\times d}$, $x\in{\mathbf{R}}^{d}$, and $b\in{\mathbf{R}}^{n}$ with $n<d$. We denote the optimal solution as $x^{\star}$. To solve (1), we apply projected gradient descent with step size $1/L$, where $L$ is the maximum eigenvalue of $A^{T}A$, obtaining iterations of the form
$x^{k+1}=\Pi_{[0,1]}\big(x^{k}-(1/L)\nabla f(x^{k})\big)=\Pi_{[0,1]}\big(x^{k}-(1/L)A^{T}(Ax^{k}-b)\big),\quad k=0,1,\dots,$
where $\Pi_{[0,1]}$ is an elementwise projection onto the interval $[0,1]$. It is well-known [5, 54], that, for this algorithm, the worst-case convergence behavior in terms of suboptimality is
$f(x^{K})-f(x^{\star})\leq\frac{L}{2K}\|x^{0}-x^{\star}\|^{2}.$
However, if we specifically consider the problem instances of the form (1) with varying $b$, the actual performance can be significantly better. To see this, we construct the objective matrix $A=U\Lambda V$, with random orthogonal matrices $U\in{\mathbf{R}}^{n\times n}$ and $V\in{\mathbf{R}}^{d\times d}$, along with rectangular matrix $\Lambda\in{\mathbf{R}}^{n\times d}$ only with diagonal entries $\Lambda_{ii}$ for $i=1,\dots,n$, drawn from the uniform distribution $\mathcal{U}[0,\sqrt{L}]$ with $L=1$. We create $1000$ instances of (1) by sampling the corresponding vectors $b$ with i.i.d. entries $b_{i}\sim\mathcal{N}(0,1)$.
Figure 1: Convergence bound of projected gradient descent on the box-constrained least-squares problem (1). The theoretical $O(K^{-1})$ worst-case bound on $f(x^{K})-f^{\star}$ is pessimistic when compared to the sample 90th-percentile and sample average.
Figure 1 shows the function-value gap across iterations for different bounds: the theoretical worst-case, the empirical average, and the 90th-percentile across instances. This highlights a commonly observed discrepancy between worst-case convergence guarantees and actual performance observed in practice.
1.1 Contributions
In this paper, we develop a data-driven performance estimation framework for analyzing the convergence of deterministic first-order methods over a distribution of convex optimization problems. By combining a priori information on the function classes considered, and a posteriori information from observing the algorithm trajectories, our data-driven technique narrows the gap between theoretical and empirical performance. Our contributions are as follows:
  • We combine the performance estimation framework (PEP) with data-driven Wasserstein distributionally robust optimization (DRO) to analyze performance of a first-order method over a distribution of convex optimization problems. To achieve this, we directly formulate the analysis as a tractable conic program that depends on the observed trajectories of the algorithm on a finite set of problem instances.
  • We establish finite-sample performance guarantees on unseen problem instances in terms of the average value, conditional value-at-risk, or worst-case value on optimality measures such as function-value gap and distance-to-solution.
  • Numerical experiments on quadratic minimization, logistic regression, and Lasso show that our approach significantly reduces the pessimism of state-of-the-art computer-assisted worst-case bounds. Specifically, our bounds capture convergence behaviors that the worst-case analysis is not able to guarantee, such as point convergence for smooth convex functions. In addition, we provide numerical evidence to conjecture faster probabilistic rates of convergence by fitting the bounds obtained by our procedure.
1.2 Related works
Classical convergence analysis and algorithm design
First-order methods have been widely adopted to solve large-scale optimization problems. Unfortunately, they can require a large number of iterations to reach high-quality solutions, especially for badly-scaled problems. This led to a significant amount of effort to analyze the worst-case performance of first-order methods under different assumptions on the problem classes involved.
The convergence behavior depends critically on problem structure. Gradient descent exhibit fast linear convergence [43] or even super-linear convergence [71] when the objective function is strongly convex or satisfies error bound conditions. Without strong convexity, it converges sublinearly [11]. Similar sublinear rates hold for other methods, including the proximal point method [29, 30], the iterative shrinkage-thresholding algorithm (ISTA), primal-dual hybrid gradient (PDHG) [12], and the alternating direction method of multipliers (ADMM) when applied to non-smooth convex optimization.
To mitigate this slow convergence behavior, Nesterov introduced the famous fast gradient method (FGM) [48] which employs an auxiliary iterate sequence to improve worst-case convergence rates. Several extensions build on this concept: the optimized gradient method (OGM) for smooth convex minimization [35], the fast iterative shrinkage-thresholding algorithm (FISTA) for composite optimization [5], the accelerated proximal point method for monotone inclusion, the Halpern iteration for fixed-point problems [34, 44], and the extra anchored gradient methods for minimax optimization [79]. Most of these acceleration schemes are optimal, in the sense that their convergence rate matches the fastest achievable rate among first-order methods [47].
Nevertheless, convergence behavior of worst-case optimal methods in practice may be significantly worse when considering a specific distribution of problem instances. Non-accelerated algorithms can outperform accelerated variants [55] on certain problem examples, suggesting that worst-case optimality does not always translate to practical efficiency. To remedy this, we aim to develop a probabilistic performance analysis framework that takes into account a specific distribution of optimization problems.
Computer-assisted performance analysis
Several recent approaches in worst-case analysis use numerical solvers, by formulating an optimization problem whose objective is the worst-case convergence rate over a specific problem class. The integral-quadratic constraint (IQC) framework uses control theory to find convergence rates for strongly-convex minimization [42, 80], using semidefinite programming (SDP) techniques. For non-asymptotic convergence rates, performance estimation problem (PEP) also uses SDP tools to formulate an optimization problem whose optimal value gives finite-step worst-case bounds [21, 70]. The PEP framework can also be used to search for optimal step sizes, leading to the discovery of the worst-case optimal first-order methods including OGM [35] and OGM-G [36] for smooth convex minimization, OptISTA [32] for composite minimization, and APPM for monotone inclusions [34, 44]. For parametric linear and convex quadratic problems, we can formulate a mixed-integer linear program (MILP) to exactly calculate the worst-case convergence bounds for first-order methods [58, 57], called the verification problem (VP). Although these approaches use numerical solvers and obtain non-analytical numerical values as results, Lyapunov analysis enables the reconstruction of human-interpretable proofs for certain problem classes [3, 73]. Unfortunately, these computer-assisted techniques provide bounds only for worst-case scenarios. In this paper, we extend the PEP framework to automatically construct probabilistic performance guarantees of first-order methods.
Probabilistic convergence analysis for deterministic algorithms
Several works move away from classical worst-case analysis to study the average-case performance of an algorithm on a certain distribution of optimization problems. Notable examples include smoothed analysis for the simplex method solving linear programs under spherically-symmetric distributions [65, 7, 66] and sorting algorithms under uniform distribution [38, 33, 75]. More recent works study the average-case performance of first-order methods solving regression tasks with Gaussian-distributed features. Such setups include applying gradient descent for strongly convex least-squares problems [56, 63, 53], smooth convex quadratic optimization [16], accelerated gradient methods for bilinear optimization [19] and Hamiltonian gradient descent for distributed optimization [50]. Other than the average-cases analysis, the high-probability convergence bounds with Gaussian distributions have been established for PDHG applied to linear programs (LPs) [78]. Probabilistic approaches have also been combined with learning-to-optimize frameworks [15], obtaining probabilistic guarantees analogous to generalization bounds in statistical learning theory [61, 62, 31]. Importantly, most of these probabilistic convergence results consider performance metrics for which we already know the worst-case convergence. For example, convergence in function-value gap or norm of the gradient for smooth convex minimization. However, in practice, we may observe stronger convergence results, such as point convergence, for specific distributions of problem instances. In addition, most of this literature focuses on either average-case or high-probability convergence, without providing systematic ways to quantify such bounds for specific quantiles.
In this paper, we provide such a systematic framework, estimating the quantile information by conditional value-at-risk ($\mathbf{CVaR}$) of selected performance metric. Moreover, we show that our approach can capture the phenomenon which is yet to be proven, namely the non-asymptotic rate to point convergence of algorithm iterates.
Distributionally robust optimization (DRO)
Distributionally robust optimization (DRO) is a popular approach to model optimization problems affected by random uncertainty, without having full knowledge on the underlying distribution. DRO constructs an ambiguity set, i.e., the set of distributions containing the true distribution with high probability, using only partial information about the uncertainty [13, 40]. Common ambiguity sets are based on either a priori information, such as moments (e.g., Chebyshev ambiguity set, Gelbrich ambiguity set, Markov ambiguity set) [18, 26, 81], or a posteriori information, such as sample realizations. The latter type of ambiguity set corresponds to a ball around the empirical distribution of samples, with respect to the KL-divergence or the Wasserstein distance [77, 46]. By formulating a robust optimization problem over this ambiguity set, DRO computes a solution and a worst-case bound on the true objective, that holds with high confidence. In this paper, we construct a PEP formulation to analyze the performance of deterministic first-order methods applied to solving a distribution of optimization problems, using data-driven Wasserstein DRO. Specifically, we model the randomness in the problem instance realizations as uncertainty in the PEP problem formulation. Thanks to the flexibility of Wasserstein DRO [39, 24], we obtain bounds on the true probabilistic performance without requiring full knowledge of the underlying distribution of problem instances, but only sample realizations of the algorithm iterate trajectories.
2 Stochastic analysis of first-order methods
Consider a minimization problem
$\begin{array}[t]{ll}\textnormal{minimize}&f(x),\end{array}$
of a proper, lower-semicontinuous, and convex objective function $f\colon{\mathbf{R}}^{d}\to{\mathbf{R}}\cup\{\infty\}$, over decision variables $x\in{\mathbf{R}}^{d}$. We assume that this problem always has an optimal solution $x^{\star}\in{\mathbf{R}}^{d}$ with optimal value $f^{\star}=f(x^{\star})$ and (sub)gradient $g^{\star}=0\in\partial f(x^{\star})$.
Algorithm
Given an initial iterate $x^{0}\in{\mathbf{R}}^{d}$, we refer to the mapping $\mathcal{A}\colon(f,x^{0})\mapsto\{x^{k}\}_{k=0,1,\dots}$ as an algorithm. We say that $\mathcal{A}$ is a first-order method if the following linear span condition holds true:
$x^{k}=x^{0}+\mathop{\bf span}\{g^{0},g^{1},\dots,g^{k}\},\quad k=1,2,\dots,$
with subgradients $g^{i}\in\partial f(x^{i})$ for $i=0,\dots,k$. Note that the last subgradient $g^{k}$ appears in proximal algorithms to implicitly define $x^{k}$ [54].
Performance metric
We measure the performance of the algorithm $\mathcal{A}$ applied on $f$ from the initial iterate $x^{0}$, using a scalar-valued function $\phi^{K}(f,x^{0})$ that quantifies the behavior of the iterates $\{x^{k}\}_{k=0}^{K}$ up to iteration $K$. See Table 1 for some examples of performance metrics.
Table 1: List of performance metrics $\phi^{K}(f,x^{0})$.
performance metric $\phi^{K}(f,x^{0})$
distance-to-solution $\big\|x^{K}-x^{\star}\big\|^{2}$
squared gradient norm $\big\|\nabla f(x^{K})\big\|^{2}$
function-value suboptimality $f(x^{K})-f^{\star}$
We assume to be in a large-scale setting where underlying problem dimension $d$ is substantially larger than $K$. This assumption not only models large-scale optimization problems, but also accounts for the strong duality argument crucial for the tightness of the worst-case analysis in the PEP framework [70, 49, 20].
Function class
We refer to a collection $\mathcal{F}$ of objective functions $f$ with shared characteristics as a function class, with examples in Table 2.
Table 2: List of function classes $\mathcal{F}$, where $M\geq 0$ and $0\leq\mu\leq L\leq\infty$.
$\mathcal{F}$ function class definition
$\mathcal{C}_{M}$ convex $M$-Lipschitz functions
$\mathcal{F}_{0,\infty}$ closed, convex, and proper functions
$\mathcal{F}_{\mu,L}$ $L$-smooth $\mu$-strongly convex functions
$\mathcal{Q}_{\mu,L}$ convex quadratic functions $x\mapsto(1/2)x^{T}Qx$ with $\mu I\preceq Q\preceq LI$
Note that the worst-case convergence rate of the algorithm refers to the tightest upper bound of the performance metric $\phi^{K}$ over the function class $\mathcal{F}$. In the following example, we present the known worst-case convergence behavior of gradient descent applied to smooth convex minimization.
Running example (worst-case convergence rate of gradient descent).
Consider an $L$-smooth convex function $f\in\mathcal{F}_{0,L}$ and the corresponding gradient descent update
$x^{k+1}=x^{k}-\eta\nabla f(x^{k}),\qquad k=0,1,\dots,$
with step size $\eta>0$. When $\eta=1/L$, for any choice of $f\in\mathcal{F}_{0,L}$, initial iterate $x^{0}\in{\mathbf{R}}^{d}$, and iteration number $K$, the following bound holds true and is known to be tight [21]:
$f(x^{K})-f^{\star}\leq\frac{L}{2(2K+1)}\|{x^{0}-x^{\star}}\|^{2}.$
Initial condition
In order to meaningfully bound the performance metric $\phi^{K}(f,x^{0})$, we restrict the initial iterate $x^{0}$ to the set $\mathcal{X}(f)\subseteq{\mathbf{R}}^{d}$. See Table 3 for common examples of $\mathcal{X}(f)$.
Table 3: List of initial condition sets $\mathcal{X}(f)$, where $r\geq 0$.
initial condition $x^{0}\in\mathcal{X}(f)$
distance-to-solution $\big\|x^{0}-x^{\star}\big\|^{2}\leq r^{2}$
squared gradient norm $\big\|\nabla f(x^{0})\big\|^{2}\leq r^{2}$
function-value suboptimality $f(x^{0})-f^{\star}\leq r$
We refer to the set of all such pairs $(f,x^{0})$ as $\mathcal{H}=\{(f,x^{0})\in\mathcal{F}\times{\mathbf{R}}^{d}\mid x^{0}\in\mathcal{X}(f)\}$.
2.1 Probabilistic performance analysis
Let $\mathcal{M}(\mathcal{H})$ be the set of all probability distributions supported on $\mathcal{H}$. We model the distribution of the optimization problem instances and the initial iterates as $(f,x^{0})\sim{\mathbf{P}}$}, where ${\mathbf{P}}\in\mathcal{M}(\mathcal{H})$. To derive probabilistic convergence guarantees with respect to $\mathbf{P}$}, we define risk measures of $\phi^{K}(f,x^{0})$ as follows.
The Value-at-Risk ($\mathbf{VaR}$) with quantile parameter $\alpha\in(0,1]$ is defined as
$\mathbf{VaR}_{\alpha}(\phi^{K}(f,x^{0}))=\mathop{\rm argmin}\left\{\tau\in{\mathbf{R}}\mid{\mathbf{P}}(\phi^{K}(f,x^{0})\leq\tau)\geq 1-\alpha\right\},$
and is exactly the $(1-\alpha)$-quantile of $\phi^{K}(f,x^{0})$ over $(f,x^{0})\sim{\mathbf{P}}$. Unfortunately, computing the $\mathbf{VaR}$ is intractable [74]. Instead, a tractable approximation of the $\mathbf{VaR}$ is the conditional Value-at-Risk ($\mathbf{CVaR}$), defined as
$\mathbf{CVaR}_{\alpha}(\phi^{K}(f,x^{0}))=\inf_{t\in{\mathbf{R}}}\,\left(t+\alpha^{-1}\mathop{\bf E{}}_{(f,x^{0})\sim{\mathbf{P}}}(\phi^{K}(f,x^{0})-t)_{+}\right),$
where $(z)_{+}=\max\{z,0\}$. According to [64, Theorem 6.2], the $\mathbf{CVaR}_{\alpha}$ can be interpreted as the conditional expectation of $\phi^{K}(f,x^{0})$ exceeding its $(1-\alpha)$-quantile:
[…]
Consequently, $\mathbf{CVaR}_{\alpha}$ upper-bounds $\mathbf{VaR}_{\alpha}$. By varying the quantile parameter $\alpha$, we can interpolate between the average case ($\alpha=1$), i.e., $\mathbf{CVaR}_{1}(\phi^{K}(f,x^{0}))=\mathop{\bf E{}}_{(f,x^{0})\sim{\mathbf{P}}}(\phi^{K}(f,x^{0}))$, and the worst case ($\alpha\to 0$), i.e., $\lim_{\alpha\to 0}\mathbf{CVaR}_{\alpha}(\phi^{K}(f,x^{0}))=\sup_{(f,x^{0})\in\mathcal{H}}\,\phi^{K}(f,x^{0})$ [74, 64]. We measure the performance of first-order methods in terms of these risk measures, which are summarized in Table 4.
Table 4: List of performance measures with respect to $(f,x^{0})\sim{\mathbf{P}}$. All measures can be expressed as a function of $\mathbf{CVaR}_{\alpha}(\phi^{K}(f,x^{0}))$ with specific values of the quantile parameter $\alpha\in(0,1]$.
performance measure quantile parameter $\alpha$
average case $\mathop{\bf E{}}_{(f,x^{0})\sim{\mathbf{P}}}\left(\phi^{K}(f,x^{0})\right)$ $\alpha=1$
$\mathbf{CVaR}$ $\mathbf{CVaR}_{\alpha}(\phi^{K}(f,x^{0}))$ $\alpha\in(0,1)$
worst case $\sup_{(f,x^{0})\in\mathcal{H}}\,\phi^{K}(f,x^{0})$ $\alpha\to 0$
Evaluating these risk measures requires solving the minimization problem
$\inf_{\ell\in\mathcal{L}}\,\mathop{\bf E{}}_{(f,x^{0})\sim{\mathbf{P}}}\,\left(\ell(f,x^{0})\right),$ (2)
over the loss function $\ell\colon\mathcal{H}\to{\mathbf{R}}$. Specifically, we set $\mathcal{L}=\{\phi^{K}(f,x^{0})\}$ for the average case and $\mathcal{L}=\{t+\alpha^{-1}(\phi^{K}(f,x^{0})-t)_{+}\mid t\in{\mathbf{R}}\}$ for the $\mathbf{CVaR}$ case. Unfortunately, solving (2) presents two main challenges. First, it requires representing functions in an infinite-dimensional space. Second, it involves evaluating $\mathop{\bf E{}}_{(f,x^{0})\sim{\mathbf{P}}}$, which requires full information of $\mathbf{P}$}, rarely available in practice. In the next section, we address these issues by constructing a tractable convex optimization problem to upper-bound (2) and quantifying its level of conservatism.
3 Convex formulation of the data-driven performance estimation problem
3.1 Finite-dimensional function representation via interpolation conditions
We obtain finite-dimensional representations of $(f,x^{0})\in\mathcal{H}$ by evaluating the function $f$ and its (sub)gradient only at the points encountered by the algorithm. This is the only function information required to compute the performance metrics in Table 1.
Suppose we are given a sequence of vectors $z^{k}=(x^{k},g^{k},f^{k})\in{\mathbf{R}}^{2d+1}$ for $k\in I$, where $I$ is an arbitrary index set. We say that $\{z^{k}\}_{k\in I}$ is $\mathcal{F}$-interpolable if there exists a function $f\in\mathcal{F}$ such that
$f^{k}=f(x^{k}),\qquad g^{k}\in\partial f(x^{k}),\qquad\forall k\in I.$
For each of the function classes in Table 2, their interpolability corresponds to a set of necessary and sufficient quadratic inequalities in the entries of the tuples $\{z^{k}\}_{k\in I}$, often referred to as the interpolation conditions [21, 69, 70, 59]. In general, these inequalities are nonconvex, as shown in the following example.
Running example (Interpolation condition of $\mathcal{F}_{0,L}$).
A set of tuples $\{z^{k}\}_{k\in I}$ is $\mathcal{F}_{0,L}$-interpolable if and only if the following inequality holds for any pair $(i,j)\in I\times I$ [70, Theorem 4]:
$f^{j}-f^{i}+(g^{j})^{T}(x^{i}-x^{j})+\frac{1}{2L}\big\|g^{i}-g^{j}\big\|^{2}\leq 0.$
This is a nonconvex quadratic inequality, due to the bilinear term $(g^{j})^{T}(x^{i}-x^{j})$.
In this paper, we consider tuples coming from the algorithm iterations and optimality conditions, which correspond to the index set $I=\{\star,0,1,\dots,K\}$, where the $\star$ index corresponds to the tuple at optimality. Note that, without loss of generality, we can set $g^{\star}=0$ and $f^{\star}=0$, because of translation invariance of the function classes considered [70].
Semidefinite variable lifting
Adopting the strategy used in the PEP framework [21, 69, 70], we lift the algorithm iterates to a higher dimensional space where the interpolation conditions can be represented as a set of convex constraints. Given a set of tuples $\{z^{k}\}_{k\in I}$ with $z^{k}=(x^{k},g^{k},f^{k})$, define the symmetric positive semidefinite Gram matrix $G\in{\mathbf{S}}_{+}^{K+2}$ and the function-value vector $F\in{\mathbf{R}}^{K+1}$ as
$G=P^{T}P,\qquad F=\left(f^{0}-f^{\star},\dots,f^{K}-f^{\star}\right),$ (3)
where
$P=\begin{bmatrix}|&|&&|\\ (x^{0}-x^{\star})&g^{0}&\cdots&g^{K}\\ |&|&&|\end{bmatrix}\in{\mathbf{R}}^{d\times(K+2)}.$ (4)
It is well-known that the interpolation conditions for the function classes in Table 2 can be written as a set of linear matrix inequalities (LMIs) in terms of $G$ and $F$ [70]. For simplicity, we write such inequalities in terms of the inner product $\big\langle\cdot,\cdot\big\rangle$ defined as
$\big\langle(X,Y),(G,F)\big\rangle=\mathop{\bf tr}(X^{T}G)+Y^{T}F,$
for any $(X,Y),\,(G,F)\in{\mathbf{S}}^{K+2}\times{\mathbf{R}}^{K+1}$. Similarly, the performance metrics in Table 1 and the initial conditions in Table 3 are affine expressions in $(G,F)$, as shown in the following example.
Running example (Interpolation conditions as LMIs).
Consider a single update of gradient descent $x^{1}=x^{0}-\eta\nabla f(x^{0})$ where $f\in\mathcal{F}_{0,L}$. Then, the interpolation conditions can be written as the following LMIs:
$-(1/2)(G_{1,2}+G_{2,1})+(1/2L)G_{1,1}+F_{1}$ $\leq 0,$ $m$ $=(\star,0)$
$(1/2)(-G_{1,3}-G_{3,1}+\eta G_{2,3}+\eta G_{3,2})+F_{2}$ $\leq 0,$ $m$ $=(\star,1)$
$(1/2L)G_{1,1}-F_{1}$ $\leq 0,$ $m$ $=(0,\star)$
$(1/2L)G_{2,2}-F_{2}$ $\leq 0,$ $m$ $=(1,\star)$
$(\eta/2)(G_{2,3}+G_{3,2})+(1/2L)(G_{1,1}-G_{1,2}-G_{2,1}+G_{2,2})-F_{1}+F_{2}$ $\leq 0,$ $m$ $=(0,1)$
$-\eta G_{1,1}+(1/2L)(G_{1,1}-G_{1,2}-G_{2,1}+G_{2,2})+F_{1}-F_{2}$ $\leq 0,$ $m$ $=(1,0).$
These are all linear in $(G,F)$ with respect to the inner product over ${\mathbf{S}}^{K+2}\times{\mathbf{R}}^{K+1}$. For instance, the case of $m=(1,0)$ writes as
$\big\langle\left((\eta/2)(e_{2}e_{3}^{T}+e_{3}e_{2}^{T})+(1/2L)(e_{1}-e_{2})(e_{1}-e_{2})^{T},\,-e_{1}+e_{2}\right),(G,F)\big\rangle\leq 0,$
with $e_{i}$ being a unit vector whose entries are all 0 except for its $i$-th element.
For the performance metric, the function-value gap $f(x^{1})-f^{\star}$ corresponds to
$f(x^{1})-f^{\star}=\big\langle(0,e_{2}),(G,F)\big\rangle.$
Similarly, the distance-to-optimality $\|x^{1}-x^{\star}\|^{2}$ and the squared gradient norm $\|\nabla f(x^{1})\|^{2}$ correspond to
$\|x^{1}-x^{\star}\|^{2}$ $=\big\langle\big((e_{1}-(1/L)e_{2})(e_{1}-(1/L)e_{2})^{T},0\big),(G,F)\big\rangle,$
$\|\nabla f(x^{1})\|^{2}$ $=\big\langle(e_{3}e_{3}^{T},0),(G,F)\big\rangle.$
Lastly, the initial condition $\|x^{0}-x^{\star}\|\leq r$ corresponds to $\langle(e_{1}e_{1}^{T},0),\,(G,F)\rangle-r^{2}\leq 0$.
From such observations, we impose affine structure on interpolation conditions, performance metrics, and initial conditions.
Assumption 1.
The interpolation conditions of the function class $\mathcal{F}$ is a set of linear matrix inequalities of the form
$\big\langle(A_{m},b_{m}),(G,F)\big\rangle\leq 0,\qquad m\in M=\{(i,j)\in I\times I\mid i\neq j\},$
where $A_{m}\in{\mathbf{S}}^{K+2}$ and $b_{m}\in{\mathbf{R}}^{K+1}$. We write
[…]
where the linear operator $\mathcal{S}\colon{\mathbf{S}}^{K+2}\times{\mathbf{R}}^{K+1}\to{\mathbf{R}}^{|M|}$ and its adjoint operator $\mathcal{S}^{*}\colon{\mathbf{R}}^{|M|}\to{\mathbf{S}}^{K+2}\times{\mathbf{R}}^{K+1}$ are defined as
$\mathcal{S}(G,F)=\Big(-\big\langle(A_{m},b_{m}),(G,F)\big\rangle\Big)_{m\in M}\,,\qquad\mathcal{S}^{*}(y)=-\sum_{m\in M}y_{m}(A_{m},b_{m}).$
The performance metric $\phi^{K}(G,F)$ has the form
$\phi^{K}(G,F)=\big\langle(A_{\mathrm{obj}},b_{\mathrm{obj}}),(G,F)\big\rangle,$
where $A_{\mathrm{obj}}\in{\mathbf{S}}^{K+2}$ and $b_{\mathrm{obj}}\in{\mathbf{R}}^{K+1}$. Lastly, the initial condition $x^{0}\in\mathcal{X}(f)$ for $f\in\mathcal{F}$ has the form $\big\langle(A_{0},b_{0}),(G,F)\big\rangle+c_{0}\leq 0$ where $A_{0}\in{\mathbf{S}}^{K+2}$, $b_{0}\in{\mathbf{R}}^{K+1}$, and $c_{0}\in{\mathbf{R}}$. We write it as
[…]
omitting the dependence on $f$ because in the definition of $G=P^{T}P$ in (3) and (4) we center the first column of $P$ around $x^{\star}$.
Note that through transformation $(f,x^{0})\mapsto(G,F)$, we lose the information on the dimension $d$ of the underlying problem. Nevertheless, as long as $d$ is large enough, the following lemma states that the finite-dimensional representation $(G,F)$ is enough to express the behavior of the given algorithm on $(f,x^{0})$.
Lemma 1.
Assume $d\geq K+2$. Under Assumption 1, the following statements hold:
  • (i)
    Given $(f,x^{0})\in\mathcal{H}$, let $\{z^{k}\}_{k\in I}$ be the set of tuples generated by the algorithm $\mathcal{A}$. Then, $(G,F)$ defined as in (3) is an element of $\mathcal{F}\cap\mathcal{X}$.
  • (ii)
    Given $(G,F)\in\mathcal{F}\cap\mathcal{X}$, there exists $(f,x^{0})\in\mathcal{H}$ such that $x^{0}\in{\mathbf{R}}^{d}$ and $P$ defined as in (4) satisfies the equation (3).
Proof.
(ii) immediately follows from $\mathcal{F}$-interpolability. For (iiii), since $G\in{\mathbf{S}}_{+}^{K+2}$ is a positive semidefinite matrix, it admits a decomposition $G=P^{T}P$ where $P\in{\mathbf{R}}^{d\times(K+2)}$. Let $x^{\star}=0$, $g^{\star}=0$, $f^{\star}=0$, and vectors $x^{0},g^{0},\dots,g^{K}$ correspond to each column of $P$. Finally, let $f^{k}$ be $(k+1)$-th entry of $F$ for $k=0,\dots,K$. Then $\{z^{k}\}_{k\in I}=\{(x^{k},g^{k},f^{k})\}_{k\in I}$ is $\mathcal{F}$-interpolable, so there exists $f\in\mathcal{F}$ and $x^{0}\in{\mathbf{R}}^{d}$ such that (3) and (4) hold true. ∎
From Gram matrix distances to function distances
We now discuss how the distance between functions and the distance between lifted variable representations are related.
For notational simplicity, let $Z=(G,F)$ be the pair of Gram matrix and function-value vector. Define the distance between two proper, lower-semicontinuous, and convex functions $f_{1}$ and $f_{2}$ with compact domain $\mathcal{D}\subseteq{\mathbf{R}}^{d}$ as
$\big\|f_{1}-f_{2}\big\|=\sup_{x\in\mathcal{D}}\,|f_{1}(x)-f_{2}(x)|+\sup_{x\in\mathcal{D}}\,\big\|\nabla f_{1}(x)-\nabla f_{2}(x)\big\|.$ (5)
In functional analysis, this defines a complete normed space called the Sobolev space. The classes of function $\mathcal{F}$ we consider are closed subspaces of such a Sobolev space, as shown in Theorem 4 in Appendix A, and therefore complete normed spaces. Define the distance over $\mathcal{H}$ as
$\|(f_{1},x_{1}^{0})-(f_{2},x_{2}^{0})\|=\|f_{1}-f_{2}\|+\|x_{1}^{0}-x_{2}^{0}\|,$
for any $(f_{1},x_{1}^{0})$ and $(f_{2},x_{2}^{0})$ in $\mathcal{H}$, and the distance over $\mathcal{F}\cap\mathcal{X}$ as
$\big\|Z_{1}-Z_{2}\big\|=\sqrt{\big\|G_{1}-G_{2}\big\|_{F}^{2}+\big\|F_{1}-F_{2}\big\|^{2}},$ (6)
for any $Z_{1}=(G_{1},F_{1})$ and $Z_{2}=(G_{2},F_{2})$ in $\mathcal{F}\cap\mathcal{X}$, where $\big\|\cdot\big\|_{F}$ is the Frobenius norm.
We claim that, according to these distances, whenever the pairs of function and initial iterate are close to each other, the corresponding Gram matrices and function-value vectors are close as well. See Appendix A for the proof.
Theorem 1.
Suppose $L=\sup_{(f,x)\in\mathcal{F}\times\mathcal{D}}\textnormal{Lip}(\nabla f)<\infty$. Given the iteration number $K$, there exists $C_{K}>0$ such that
$\big\|Z_{1}-Z_{2}\big\|\leq C_{K}\|(f_{1},x_{1}^{0})-(f_{2},x_{2}^{0})\|,$
for any $x_{1}^{0},x_{2}^{0}\in\mathcal{D}$ and any closed, convex, and proper functions $f_{1},f_{2}\in\mathcal{F}$ with $f_{1}^{\star}=f_{2}^{\star}=0$, $g_{1}^{\star}=g_{2}^{\star}=0$, and $x_{1}^{\star}=x_{2}^{\star}=0$.
Remark 1.
Assumptions of Theorem 1 are true for function classes with Lipschitz gradients, e.g., $\mathcal{F}=\mathcal{F}_{0,L}$, and for bounded initial iterates.
3.2 Approximating the expectation from data
After the variable lifting from Section 3.1, the expectation in (2) equals
$\mathop{\bf E{}}_{Z\sim{\mathbf{P}}_{Z}}\left(\ell(Z)\right),$ (7)
where $Z=(G,F)$ and ${\mathbf{P}}_{Z}\in\mathcal{M}(\mathcal{F}\cap\mathcal{X})$ is a probability distribution induced from ${\mathbf{P}}\in\mathcal{M}(\mathcal{H})$ by mapping $(f,x^{0})\mapsto Z$. Unfortunately, in most practical scenarios, complete knowledge of ${\mathbf{P}}_{Z}$ is unavailable. We address this issue using distributionally-robust optimization (DRO) [46, 39, 40]. Specifically, we incorporate a priori structural information modeled in the previous sections, such as initial conditions and the interpolation conditions, with data-driven a posteriori information from running the algorithm on sample realizations of $(f,x^{0})\sim{\mathbf{P}}$.
Distributionally-robust optimization
Consider a set $\mathcal{P}$ of probability distributions supported on $\mathcal{F}\cap\mathcal{X}$. If ${\mathbf{P}}_{Z}\in\mathcal{P}$, then the worst-case expectation over $\mathcal{P}$ upper-bounds the true expectation (7) over ${\mathbf{P}}_{Z}$, i.e.,
$\inf_{\ell\in\mathcal{L}}\,\mathop{\bf E{}}_{(f,x^{0})\sim{\mathbf{P}}}\left(\ell(f,x^{0})\right)=\inf_{\ell\in\mathcal{L}}\,\underset{Z\sim{\mathbf{P}}_{Z}}{\mathop{\bf E{}}}\left(\ell(Z)\right)\leq\begin{array}[t]{ll}\underset{\mathbf{Q}_{Z}\in\mathcal{P}}{\sup}&\underset{\ell\in\mathcal{L}}{\inf}\,\underset{Z\sim\mathbf{Q}_{Z}}{\mathop{\bf E{}}}\left(\ell(Z)\right).\end{array}$ (8)
Our goal is to construct an ambiguity set $\mathcal{P}$ that is guaranteed to contain ${\mathbf{P}}_{Z}$ with high probability, using only partial information about ${\mathbf{P}}_{Z}$ obtained from samples $\widehat{Z}_{1},\dots,\widehat{Z}_{N}\sim{\mathbf{P}}_{Z}$.
Finite-sample guarantee of Wasserstein DRO
Consider the empirical distribution $\widehat{{\mathbf{P}}}_{Z}=(1/N)\sum_{i=1}^{N}\delta_{\widehat{Z}_{i}}$, where $\delta_{\widehat{Z}_{i}}$ is the Dirac distribution at $i$-th sample $\widehat{Z}_{i}=(\widehat{G}_{i},\widehat{F}_{i})$. We set the ambiguity set $\mathcal{P}$ as a ball of radius $\varepsilon$ around $\widehat{{\mathbf{P}}}_{Z}$ defined as
[…]
with respect to the type-$1$ Wasserstein distance
$W({\mathbf{P}}_{Z},\mathbf{Q}_{Z})$ $=\inf_{\pi\in\Pi({\mathbf{P}}_{Z},\mathbf{Q}_{Z})}\,\left(\int_{\mathcal{F}\cap\mathcal{X}}\big\|Z_{1}-Z_{2}\big\|\,\pi(dZ_{1},dZ_{2})\right),$
where $\|\cdot\|$ is a norm over $\mathcal{F}\cap\mathcal{X}$ and $\pi$ is a joint distribution of $Z_{1}$ and $Z_{2}$ with marginals ${\mathbf{P}}_{Z}$ and $\mathbf{Q}_{Z}$.
Assume that the true distribution ${\mathbf{P}}_{Z}$ is light-tailed, i.e., $\mathop{\bf E{}}_{Z\sim{\mathbf{P}}_{Z}}\,[\exp\left(\big\|Z\big\|^{a}\right)]<\infty$ for some $a>1$. Given a confidence level $\beta\in(0,1)$, choose a radius $\varepsilon=\varepsilon_{N}(\beta)>0$ as
$\varepsilon_{N}(\beta)=\begin{cases}\left(\frac{\log(c_{1}\beta^{-1})}{c_{2}N}\right)^{\left((K+2)(K+3)/2+(K+1)\right)^{-1}}&\textnormal{ if }N\geq\frac{\log(c_{1}\beta^{-1})}{c_{2}}\\ \left(\frac{\log(c_{1}\beta^{-1})}{c_{2}N}\right)^{1/a}&\textnormal{ if }N<\frac{\log(c_{1}\beta^{-1})}{c_{2}},\end{cases}$ (9)
with some constants $c_{1},c_{2}>0$. Then, from well-known measure concentration result [22, Theorem 2], the true distribution ${\mathbf{P}}_{Z}$ lies in $B_{{\varepsilon}}$ with probability at least $1-\beta$, i.e.,
${\mathbf{P}}_{Z}^{N}\left(\widehat{{\mathbf{P}}}_{Z}\mid W({\mathbf{P}}_{Z},\widehat{{\mathbf{P}}}_{Z})\leq\varepsilon\right)\geq 1-\beta.$ (10)
Consequently, the worst-case expectation in inequality (8) upper-bounds the true expectation (7) with probability at least $1-\beta$. In the next section, we develop a tractable formulation to compute the worst-case expected loss.
Connection to worst-case expectation over infinite-dimensional function spaces
For the sake of completeness, we discuss an alternative way to derive the worst-case expectation over the lifted Gram matrix formulation in (8), using the worst-case expectation over the infinite-dimensional space $\mathcal{H}$ of functions and initial iterates. Consider the relation:
$\underset{(f,x^{0})\sim{\mathbf{P}}}{\mathop{\bf E{}}}\,\big(\ell(f,x^{0})\big)$$\underset{Z\sim{\mathbf{P}}^{Z}}{\mathop{\bf E{}}}\left(\ell(Z)\right)$$\underset{\mathbf{Q}\in\mathcal{P}}{\sup}\,\underset{(f,x^{0})\sim\mathbf{Q}}{\mathop{\bf E{}}}\,\left(\ell(f,x^{0})\right)$$\underset{\mathbf{Q}^{Z}\in\mathcal{P}^{Z}}{\sup}\,\underset{Z\sim\mathbf{Q}^{Z}}{\mathop{\bf E{}}}\left(\ell(Z)\right)$}.} w.h.p $\scriptstyle$$\leq\text${ w.h.p }} w.h.p $\scriptstyle$$\leq\text${ w.h.p }}Prop. 1
Although not explicitly stated, the finite-sample guarantee also holds for distributions supported on infinite-dimensional function spaces from Table 2 [41]. We claim that the worst-case expectation over the distributions supported on the set $\mathcal{F}\cap\mathcal{X}$ upper-bounds the worst-case expectation over the distributions supported on $\mathcal{H}$.
Proposition 1.
Given $\varepsilon>0$, there exists a Wasserstein ambiguity set $\mathcal{P}\subseteq\mathcal{M}(\mathcal{H})$ such that
$\underset{\mathbf{Q}\in\mathcal{P}}{\sup}\,\underset{(f,x^{0})\sim\mathbf{Q}}{\mathop{\bf E{}}}\big(\ell(f,x^{0})\big)\leq\underset{\mathbf{Q}^{Z}\in B_{{\varepsilon}}}{\sup}\,\underset{Z\sim\mathbf{Q}^{Z}}{\mathop{\bf E{}}}\left(\ell(Z)\right).$
Proof.
See Appendix B. ∎
3.3 Tractable convex formulation using duality theory
We now derive a convex formulation of the upper bound in (8). Note that the support set $\mathcal{F}\cap\mathcal{X}$ of the distributions considered is compact, as the algorithm trajectories are constructed from compact initial iterate sets.
Assumption 2.
$\mathcal{F}\cap\mathcal{X}$ is a compact set in ${\mathbf{S}}^{K+2}\times{\mathbf{R}}^{K+1}$.
This is not only required to obtain our next result, but also implies the light-tailed property discussed in Section 3.2 which gives finite-sample guarantee (10) of Wasserstein DRO. We now present our main result, which is a convex formulation of the upper bound in (8), whose proof is deferred to Appendix C.
Theorem 2.
Given samples $(\widehat{G}_{i},\widehat{F}_{i})$ for $i=1,\dots,N$, consider the set $\mathcal{L}$ of loss functions $\ell\colon\mathcal{F}\cap\mathcal{X}\to{\mathbf{R}}$ parametrized as
$\mathcal{L}=\left\{\ell(G,F)=\max_{1\leq j\leq J}\,\ell^{j}(G,F)\;\middle|\;\ell^{j}(G,F)=\big\langle(A_{\mathrm{obj}}^{j},b_{\mathrm{obj}}^{j}),(G,F)\big\rangle+c_{\mathrm{obj}}^{j}t,~t\in{\mathbf{R}}\right\},$
with $A_{\mathrm{obj}}^{j}\in{\mathbf{S}}^{K+2}$, $b_{\mathrm{obj}}^{j}\in{\mathbf{R}}^{K+1}$, and $c_{\mathrm{obj}}^{j}\in{\mathbf{R}}$. Under Assumption 1, for every $\varepsilon>0$,
$\begin{array}[t]{lll}\underset{\mathbf{Q}_{Z}\in B_{{\varepsilon}}}{\sup}&\underset{\ell\in\mathcal{L}}{\inf}&\underset{(G,F)\sim\mathbf{Q}_{Z}}{\mathop{\bf E{}}}\,\left(\ell(G,F)\right)\end{array}=\begin{array}[t]{lll}\textnormal{maximize}&\sum_{i=1}^{N}\sum_{j=1}^{J}\alpha_{i}^{j}\big\langle(A_{\mathrm{obj}}^{j},b_{\mathrm{obj}}^{j}),\,(G_{i}^{j},F_{i}^{j})\big\rangle\\ \textnormal{subject to}&\sum_{i=1}^{N}\sum_{j=1}^{J}\alpha_{i}^{j}\|(G_{i}^{j},F_{i}^{j})-(\widehat{G}_{i},\widehat{F}_{i})\|\leq\varepsilon,\end{array}$
which is also equivalent to
$\begin{array}[t]{lll}\textnormal{minimize}&(1/N)\sum_{i=1}^{N}s_{i}\\ \textnormal{subject to}&c_{\mathrm{obj}}^{j}t-c_{0}\tau_{i}^{j}-\big\langle(X_{i}^{j},Y_{i}^{j}),\,(\widehat{G}_{i},\widehat{F}_{i})\big\rangle+\lambda\varepsilon\leq s_{i}\\ &-\mathcal{S}^{*}(y_{i}^{j})-(X_{i}^{j},Y_{i}^{j})+\tau_{i}^{j}(A_{0},b_{0})-(A_{\mathrm{obj}}^{j},b_{\mathrm{obj}}^{j})\in{\mathbf{S}}_{+}^{K+2}\times\{0\}\\ &\|(X_{i}^{j},Y_{i}^{j})\|_{*}\leq\lambda,\hfill i=1,\dots,N,\;j=1,\dots,J.\end{array}$ (DRO-PEP)
Here, the primal variables are $(\alpha_{i}^{j})\in\Gamma$ and $(G_{i}^{j},F_{i}^{j})\in\mathcal{F}\cap\mathcal{X}$ where
[…]
and dual variables are $s_{i}\in{\mathbf{R}}$, $t\in{\mathbf{R}}$, $\tau_{i}^{j}\in{\mathbf{R}}_{+}$, $(X_{i}^{j},Y_{i}^{j})\in{\mathbf{S}}^{K+2}\times{\mathbf{R}}^{K+1}$, $\lambda\in{\mathbf{R}}_{+}$, and $y_{i}^{j}\in{\mathbf{R}}_{+}^{|M|}$.
Remark 2 (Problem complexity).
The maximization problem of Theorem 2 is a finite-dimensional nonconvex conic program, whereas the minimization problem (DRO-PEP) is a convex conic program. Still, there is a trade-off between the problem complexity of (DRO-PEP), which depends on the numbers of iterations $K$ and samples $N$, and the confidence $1-\beta$ for which the inequality (8) is valid.
Remark 3 (Constructing a worst-case distribution).
According to [46, Corollary 4.6], a worst-case distribution achieving the optimal value of (DRO-PEP) is a weighted sum of Dirac distributions. Let $(\alpha_{i}^{j})\in\Gamma$ and $(G_{i}^{j},F_{i}^{j})$ for $i=1,\dots,N$ and $j=1,\dots,J$ be optimal solutions to the maximization problem of Theorem 2. For each tuple $(G_{i}^{j},F_{i}^{j})$, there exists an interpolating function $f_{i}^{j}\in\mathcal{F}$ and an initial point $(x_{i}^{j})^{0}\in{\mathbf{R}}^{d}$. These can be constructed by factorizing $G_{i}^{j}$ as $G_{i}^{j}=(P_{i}^{j})^{T}P_{i}^{j}$ and constructing the interpolating function $f_{i}^{j}\in\mathcal{F}$, as in the proof of Lemma 1. Note that there can be multiple such functions and initial points. Then a worst-case distribution is constructed as $\mathbf{Q}_{Z}^{\star}=\sum_{i=1}^{N}\sum_{j=1}^{J}\alpha_{i}^{j}\delta_{i}^{j}$, where each $\delta_{i}^{j}$ is a Dirac distribution on every $(f_{i}^{j},(x_{i}^{j})^{0})$. Evaluating the expectation of this worst-case distribution gives the optimal value of (DRO-PEP).
In the following corollary, we specialize Theorem 2 to upper-bound the true performance in terms of expectation and $\mathbf{CVaR}$.
Corollary 2.1.
Under Assumption 2, consider the setup of Theorem 2. Let the confidence $\beta\in(0,1)$ and the ambiguity set radius $\varepsilon$ be given as in (9). The following statements hold:
  • Expectation: With probability at least $1-\beta$, the average-case performance $\mathop{\bf E{}}\big(\phi^{K}(Z)\big)$ is upper-bounded by the optimal value of problem (DRO-PEP) with $\alpha=1$, $J=1$, and $(A_{\mathrm{obj}}^{1},b_{\mathrm{obj}}^{1},c_{\mathrm{obj}}^{1})=(A_{\mathrm{obj}},b_{\mathrm{obj}},0)$.
  • $\mathbf{CVaR}$: With probability at least $1-\beta$, the $\mathbf{CVaR}$-performance $\mathbf{CVaR}_{\alpha}(\phi^{K}(Z))$ with $\alpha\in(0,1]$ is upper-bounded by the optimal value of problem (DRO-PEP) with $J=2$, $(A_{\mathrm{obj}}^{1},b_{\mathrm{obj}}^{1},c_{\mathrm{obj}}^{1})=(\alpha^{-1}A_{\mathrm{obj}},\alpha^{-1}b_{\mathrm{obj}},1-\alpha^{-1})$, and $(A_{\mathrm{obj}}^{2},b_{\mathrm{obj}}^{2},c_{\mathrm{obj}}^{2})=(0,0,1)$.
Proof.
From the finite-sample guarantee (10), the true distribution ${\mathbf{P}}_{Z}$ is an element of the Wasserstein ambiguity set $B_{{\varepsilon}}$ with probability at least $1-\beta$. Therefore, $\mathop{\bf E{}}_{Z\sim{\mathbf{P}}_{Z}}(\phi^{K}(Z))$ and $\mathbf{CVaR}_{\alpha}(\phi^{K}(Z))$ of true distribution ${\mathbf{P}}_{Z}$ are all upper-bounded by the corresponding worst-case expectations over $\mathbf{Q}_{Z}\in B_{{\varepsilon}}$, with probability at least $1-\beta$. Applying Theorem 2 with the given choices of loss functions, we get reformulations of such worst-case expectations as convex optimization problems. In Appendix C, we give the complete formulations. ∎
For our running example, we provide the the convex reformulation from Corollary 2.1 for the expectation case.
Running example (Deriving (DRO-PEP) for gradient descent step).
The adjoint operator $\mathcal{S}^{*}\colon{\mathbf{R}}^{6}\to{\mathbf{S}}^{3}\times{\mathbf{R}}^{2}$ of Assumption 1 is
$\mathcal{S}^{*}(y_{i})$ $=y_{i,1}\left(-(1/2)(e_{1}e_{2}^{T}+e_{2}e_{1}^{T})+(1/2L)e_{1}e_{1}^{T},\,e_{1}\right)$
$\quad+y_{i,2}\left(-(1/2)(e_{1}e_{3}^{T}+e_{3}e_{1}^{T})+(\eta/2)(e_{2}e_{3}^{T}+e_{3}e_{2}^{T}),\,e_{2}\right)$
$\quad+y_{i,3}\left((1/2L)e_{1}e_{1}^{T},\,-e_{1}\right)$
$\quad+y_{i,4}\left((1/2L)e_{2}e_{2}^{T},\,-e_{2}\right)$
$\quad+y_{i,5}\left((\eta/2)(e_{2}e_{3}^{T}+e_{3}e_{2}^{T})+(1/2L)(e_{1}-e_{2})(e_{1}-e_{2})^{T},\,-e_{1}+e_{2}\right)$
$\quad+y_{i,6}\left(-\eta e_{1}e_{1}^{T}+(1/2L)(e_{1}-e_{2})(e_{1}-e_{2})^{T},\,e_{1}-e_{2}\right),$
where $y_{i}=(y_{i,m})_{m\in M}\in{\mathbf{R}}^{6}$. According to Corollary 2.1, we get
$\sup_{\mathbf{Q}\in B_{{\varepsilon}}}\mathop{\bf E{}}(\phi^{K}(G,F))=\begin{array}[t]{lll}\textnormal{minimize}&(1/N)\sum_{i=1}^{N}s_{i}\\ \textnormal{subject to}&\tau_{i}-\big\langle(X_{i},Y_{i}),\,(\widehat{G}_{i},\widehat{F}_{i})\big\rangle+\lambda\varepsilon\leq s_{i}\\ &-\mathcal{S}^{*}(y_{i})-(X_{i},Y_{i})+\tau_{i}(e_{1}e_{1}^{T},0)-(0,e_{2})\in{\mathbf{S}}_{+}^{3}\times\{0\}\\ &\big\|(X_{i},Y_{i})\big\|_{*}\leq\lambda,\hfill i=1,\dots,N,\end{array}$
with variables $s_{i}\in{\mathbf{R}}$, $\tau_{i}\in{\mathbf{R}}_{+}$, $(X_{i},Y_{i})\in{\mathbf{S}}^{3}\times{\mathbf{R}}^{2}$, $\lambda\in{\mathbf{R}}_{+}$, and $y_{i}\in{\mathbf{R}}_{+}^{6}$.
3.4 Bridging worst-case and average-case analyses
The radius $\varepsilon$ of Wasserstein ambiguity set $B_{{\varepsilon}}$ controls the trade-off between a priori and a posteriori information. When $\varepsilon$ gets larger, the upper bounds in Corollary 2.1 approach the worst-case convergence bound, which only uses a priori information, i.e., function classes and initial conditions. On the other hand, when $\varepsilon$ gets smaller, the bounds from Corollary 2.1 approach the a posteriori empirical performance, i.e., the sampled average-case and $\mathbf{CVaR}$-case losses for the observed iterate trajectories. The following theorem states the behavior of the problem (DRO-PEP) on its extreme cases: $\varepsilon\to 0$ and $\varepsilon\to\infty$.
Theorem 3.
Consider the setup of Corollary 2.1.
  • When $\varepsilon\to 0$, the optimal values of (DRO-PEP) converge respectively to the in-sample average-case performance $(1/N)\sum_{i=1}^{N}\phi^{K}(\widehat{Z}_{i})$ and the in-sample $\mathbf{CVaR}$-performance $\inf_{t\in{\mathbf{R}}}\big(t+(\alpha^{-1}/N){\textstyle\sum_{i=1}^{N}}(\phi^{K}(\widehat{Z}_{i})-t)_{+}\big)$ which is the optimal value of
    $\begin{array}[t]{ll}\textnormal{maximize}&(\alpha^{-1}/N)\sum_{i=1}^{N}\alpha_{i}\phi^{K}(\widehat{Z}_{i})\\ \textnormal{subject to}&(1/N)\sum_{i=1}^{N}\alpha_{i}=\alpha,\end{array}$
    with variables $\alpha_{i}\in[0,1]$.
  • Under Assumption 2, there exists $\bar{\varepsilon}>0$ such that for any $\varepsilon\geq\bar{\varepsilon}$, the optimal values of (DRO-PEP) all equal the worst-case performance:
    $\underset{Z\in\mathcal{F}\cap\mathcal{X}}{\sup}\phi^{K}(Z)=\begin{array}[t]{ll}\textnormal{minimize}&-c_{0}\tau\\ \textnormal{subject to}&-\mathcal{S}^{*}(y)+\tau(A_{0},b_{0})-(A_{\mathrm{obj}},b_{\mathrm{obj}})\in{\mathbf{S}}_{+}^{K+2}\times\{0\},\end{array}$
    with variables $\tau\in{\mathbf{R}}_{+}$ and $y\in{\mathbf{R}}_{+}^{|M|}$. This is exactly the dual of worst-case PEP [70, Section 3.4].
Proof.
First, consider the case when $\varepsilon\to 0$, where the maximization problem of Theorem 2 has a unique feasible point $Z_{i}=\widehat{Z}_{i}$ for $i=1,\dots,N$. Then, the objective value of (DRO-PEP) in case of $J=1$ will be the in-sample average $(1/N)\sum_{i=1}^{N}\phi^{K}(\widehat{Z}_{i})$. Similarly, the objective value of (DRO-PEP) in case of $J=2$ converges to $\max\,\{(\alpha^{-1}/N)\sum_{i=1}^{N}\alpha_{i}\phi^{K}{(\widehat{Z}_{i})}\mid(1/N)\sum_{i=1}^{N}\alpha_{i}=\alpha,\,\alpha_{i}\in[0,1]\}$, which is exactly the $\mathbf{CVaR}_{\alpha}(\phi^{K}(Z))$ of empirical distribution $\widehat{{\mathbf{P}}}_{Z}$.
Under Assumption 2, there exists a worst-case instance $\bar{Z}\in\mathcal{F}\cap\mathcal{X}$ such that $\sup_{Z\in\mathcal{F}\cap\mathcal{X}}\phi^{K}(Z)=\phi^{K}(\bar{Z})$. Let $\bar{\varepsilon}=\max_{1\leq i\leq N}\|\bar{Z}-\widehat{Z}_{i}\|$. Then for any $\varepsilon\geq\bar{\varepsilon}$, $\bar{Z}$ is a feasible point of maximization problem in Theorem 2 since
${\textstyle\sum_{i=1}^{N}\sum_{j=1}^{J}}\alpha_{i}^{j}\|\bar{Z}-\widehat{Z}_{i}\|=(1/N){\textstyle\sum_{i=1}^{N}}\|\bar{Z}-\widehat{Z}_{i}\|\leq\bar{\varepsilon}\leq\varepsilon,$
[…]
The following example illustrates how (DRO-PEP) interpolates the probabilistic performance of empirical distribution and the worst-case convergence bound, as Theorem 3 suggests.
Running example (Gradient descent on smooth convex minimization).
Consider gradient descent iterations $x^{k+1}=x^{k}-\eta\nabla f(x^{k})$ with $f\in\mathcal{F}_{0,L}$ and $\eta\in(1/L,2/L)$, given the initial condition $\|x^{0}-x^{\star}\|\leq r$. To show the effects of (DRO-PEP) in a data-driven procedure, we fix $L=1$, $r=1$, and $\eta=3/(2L)$ and solve the first optimization problem of Corollary 2.1 for $K=1,5,10$, with 20 logarithmically spaced values $\varepsilon\in[10^{-5},10^{2}]$. Results are shown in Figure 2.
For each value of $K$, we observe that the (DRO-PEP)-objective is increasing in $\varepsilon$ until reaching the $\bar{\varepsilon}$-threshold described in Theorem 3, at which point and further $\varepsilon\geq\bar{\varepsilon}$, the (DRO-PEP)-objective matches the theoretical worst-case bound [21, 37]:
$\sup_{f\in\mathcal{F}}\,(f(x^{K})-f^{\star})\leq\frac{Lr^{2}}{2}\max\left\{\frac{1}{2LK\eta+1},\,(1-L\eta)^{2K}\right\}.$
Figure 2: Optimal objective value of average-case performance problem in Corollary 2.1 for gradient descent with $N=50$ samples of convex quadratic functions and $x^{0}=(r,0,\dots,0)\in{\mathbf{R}}^{d}$ with $d=200$. For each $K$ value, the top and bottom dashed lines give the theoretical worst-case bound and in-sample average, respectively.
3.5 Wasserstein distance with relative error
In this section, we introduce how we choose the norm over $\mathcal{F}\cap\mathcal{X}$, in order to give equal importance to each iteration along the trajectory. Specifically, the norm described in (6) overly weights earlier iterations because they tend to have larger entries, e.g., gradient norm or suboptimality gap. To address this issue, consider the scaled norm $\|\cdot\|_{D}$} over $\mathcal{F}\cap\mathcal{X}$ defined as
$\big\|(G,F)\big\|_{D}=\sqrt{D^{1/2}(G,F)D^{1/2}}=\big\|{(D_{G}^{1/2}GD_{G}^{1/2},D_{F}F)}\big\|,$
with a pair of positive definite diagonal matrices $D=(D_{G},D_{F})\in{\mathbf{S}}_{++}^{K+2}\times{\mathbf{S}}_{++}^{K+1}$ referred to as a preconditioner. An effective way to assign higher weight to later iterations is to construct a preconditioner that is inversely proportional to the observed values of the entries in the diagonal of the samples $\widehat{G}_{i}$ and the entries of $\widehat{F}_{i}$.
We use the averaged value of sampled gradient magnitude and function-value gap for every $k$-th iterate as normalizing factor for $g^{k}$ and $f^{k}-f^{\star}$, respectively. Furthermore, in order for the quantity $\big\|(G_{1},F_{1})-(G_{2},F_{2})\big\|_{D}$ to be similar for any choice of total iteration number $K$, we further normalize $D$ by the number of its diagonal components. Therefore, we use the following form of the preconditioner $D=(D_{G},D_{F})\in{\mathbf{S}}_{++}^{K+2}\times{\mathbf{S}}_{++}^{K+1}$ defined as
$D_{G}$ $=\frac{N^{2}}{K+2}\mathop{\bf diag}\left({\textstyle\sum_{i=1}^{N}}\big\|\hat{x}_{i}^{0}-\hat{x}_{i}^{\star}\big\|,\;{\textstyle\sum_{i=1}^{N}}\big\|\hat{g}_{i}^{0}\big\|,\;\cdots,\;{\textstyle\sum_{i=1}^{N}}\big\|\hat{g}_{i}^{K}\big\|\right)^{-2}$
$D_{F}$ $=\frac{N}{K+1}\mathop{\bf diag}\left({\textstyle\sum_{i=1}^{N}}(\hat{f}_{i}^{0}-\hat{f}_{i}^{\star}),\cdots,{\textstyle\sum_{i=1}^{N}}(\hat{f}_{i}^{K}-\hat{f}_{i}^{\star})\right)^{-1}.$
Corollary 3.1.
Consider the setup of Theorem 2 with the Wasserstein ambiguity set $B_{{\varepsilon}}$ defined with respect to the norm $\|\cdot\|_{D}$ over $\mathcal{F}\cap\mathcal{X}$. The convex reformulation is given by (DRO-PEP) with the dual norm $\|\cdot\|_{*}$ replaced by $\|\cdot\|_{D^{-1}}$.
Proof.
Since the dual norm $\big\|\cdot\big\|_{D*}$ of $\big\|\cdot\big\|_{D}$ is evaluated as
$\big\|(X,Y)\big\|_{D*}$ $=\sup_{\big\|(G,F)\big\|_{D}\leq 1}\big\langle(X,Y),(G,F)\big\rangle$
$=\sup_{\|D^{1/2}(G,F)D^{1/2}\|\leq 1}\big\langle(D_{G}^{-1/2}XD_{G}^{-1/2},D_{F}^{-1}Y),D^{1/2}(G,F)D^{1/2}\big\rangle$
$=\big\|\left(D_{G}^{-1/2}XD_{G}^{-1/2},\,D_{F}^{-1}Y\right)\big\|=\big\|(X,Y)\big\|_{D^{-1}},$
where $D^{-1}=(D_{G}^{-1},D_{F}^{-1})$, we can replace the dual norm constraint $\|(X_{i}^{j},Y_{i}^{j})\|_{*}\leq\lambda$ in (DRO-PEP) by $\|(X_{i}^{j},Y_{i}^{j})\|_{D^{-1}}\leq\lambda$. ∎
3.6 Data-driven convergence rates
Thus far, we have constructed numeric bounds for fixed values of iteration number $K$. We now turn into extrapolating these bounds to larger values of $K$. Specifically, we propose a method to fit curves that upper bound the numeric values obtained by solving (DRO-PEP), obtaining data-driven convergence rates.
We say an algorithm has a sublinear convergence rate if for some $\gamma>0$, the performance metric as a function of $K$ decreases as $O(K^{-\gamma})$. In contrast, we say it has a linear convergence rate if for some $\rho\in(0,1)$, the metric decreases as $O(\rho^{K})$. The following example for gradient descent on different function classes highlights the difference.
Running example (Gradient descent rates).
Consider gradient descent iterations $x^{k+1}=x^{k}-(1/L)\nabla f(x^{k})$. When $f\in\mathcal{F}_{0,L}$, the following worst-case convergence bound holds for the function-value gap [21, 37]:
$f(x^{K})-f^{\star}\leq\dfrac{L\|x^{0}-x^{\star}\|^{2}}{2(2K+1)},$
i.e. a sublinear convergence rate with $\gamma=1$. Instead, if we consider the function class $\mathcal{F}_{\mu,L}$ with $\mu>0$, then the following worst-case convergence bound holds with a different initial condition [9, Chapter 9.3]:
$f(x^{K})-f^{\star}\leq(1-\mu/L)^{K}(f(x^{0})-f^{\star}),$
i.e. a linear convergence rate with $\rho=1-\mu/L$.
To fit these curves, we use a constrained least-squares procedure. Given pairs $\{(K,\phi_{K})\}_{K=1}^{{K_{\max}}}$, where $\phi_{K}$ is the optimal value of (DRO-PEP) with iteration number $K$, our goal is to approximately fit constants $C>0$, $\gamma\geq 0$, and $\rho\in(0,1]$ to the form $\phi_{K}\approx C\rho^{K}(K+1)^{-\gamma}$. Note that we use $(K+1)$ for sublinear terms in order to eliminate the effect of $\gamma$ when $K=0$. Taking the log transform of both sides gives the fit
$\log\phi_{K}\approx\log C+K\log\rho-\gamma\log(K+1).$ (11)
Let $H\in{\mathbf{R}}^{{K_{\max}}\times 3}$ be the matrix whose entries, for every row $K$, are $H_{K1}=1$, $H_{K2}=-\log(K+1)$ and $H_{K3}=K$, and also let $h\in{\mathbf{R}}^{{K_{\max}}}$ be the vector with entries $h_{K}=\log\phi_{K}$, for $K=1,\cdots,{K_{\max}}$. Our goal is to fit (11) in a least-squares sense; however, we make two modifications. First, in order to maintain the bounds as in (8) with high confidence (10), we constrain our fitted curve to upper-bound the observed $\phi_{K}$ for each $K$. Second, since $\phi_{K}$ tends to decrease with $K$, we normalize the contribution to the least-squares objective via the diagonal matrix $W\in{\mathbf{S}}^{{K_{\max}}\times{K_{\max}}}$ with each diagonal entry taking into account the magnitude of $\phi_{K}$, by setting $W_{KK}=1+\log(\phi_{0}/\phi_{K})$. Overall, we solve the following constrained, weighted least-squares problem:
$\begin{array}[]{ll}\text{minimize}&\|W(Hx-h)\|^{2}\\ \text{subject to}&Hx\geq h.\end{array}$ (12)
From the optimal solution $x^{\star}\in{\mathbf{R}}^{3}$ to (12), we recover $\rho=e^{x^{\star}_{1}}$, $\gamma=x^{\star}_{2}$, and $C=e^{x^{\star}_{3}}$. Additionally, we may enforce either $x_{1}=0$ or $x_{2}=1$ if we decide to isolate either a sublinear or linear fit, respectively.
4 Numerical experiments
We now analyze the probabilistic performance of various first-order methods in unconstrained quadratic optimization, logistic regression, and Lasso. All code to reproduce our experiments is available at
https://github.com/stellatogrp/dro_pep.
We focus on solving the minimization problems derived in Corollary 2.1 with the modifications discussed in Section 3.5. For each experiment, we also use the curve fitting techniques from Section 3.6. All examples are written in Python 3.12 and all problems are solved with the Clarabel solver with default setting [28]. Also, we obtain the worst-case convergence bound by solving the corresponding PEP using the PEPit toolbox [27].
Parameter selection
For each experiment, we generate $N=100$ i.i.d. samples, each consisting of function realizations and corresponding iterate trajectories, to form the (DRO-PEP) problem. According to Section 3.2, the radius $\varepsilon$ of Wasserstein ambiguity set in problem (DRO-PEP) controls the likelihood that the bound (8) holds. As commonly done in the DRO literature [46, 24], use cross-validation over $\varepsilon$ values to arrive at the performance guarantee with desired confidence $1-\beta$. That is, we generate 10000 additional problems to estimate the true probability of constraint satisfaction defined in (8). The cross-validation is performed over 10 logarithmically-spaced values $\varepsilon\in[10^{-2},10^{1}]$, and we choose the smallest such $\varepsilon$ value for $\beta=0.05$. In the $\mathbf{CVaR}$ experiments, we fix $\alpha=0.1$, or the 90th-percentile.
Numerical results
For each experiment, we compare the following values across iterations:
  • Worst-case convergence bound from PEP. Using the PEPit toolbox, we compute the worst-case convergence bounds by solving the corresponding PEP.
  • DRO objective values. We solve the (DRO-PEP) problems in Corollary 2.1, to obtain the worst-case expectation and the worst-case $\mathbf{CVaR}$ of the chosen performance metric. In all plots, the solid lines show the optimal objective value of (DRO-PEP).
  • Fitted rates. We provide tables consisting of the worst-case rates and fitted convergence rates from the worst-case $\mathop{\bf E{}}(\phi^{K}(f,x^{0}))$ and $\mathbf{CVaR}_{\alpha}(\phi^{K}(f,x^{0}))$, using the approach in Section 3.6.
  • Empirical values. To simulate the true values, we also plot the empirical worst-case, average-case, and $\mathbf{CVaR}$ as computed from the 10000 out-of-sample problem instances. In all plots, the dashed lines show the empirical sample values.
  • Solve times. We provide time tables for solving (DRO-PEP), averaged over cross-validated $\varepsilon$ values, and for solving the worst-case PEP. All time tables are given in Appendix D.
4.1 Unconstrained quadratic minimization
For this experiment, we restrict ourselves to the class of $L$-smooth $\mu$-strongly convex quadratic functions, which is
$f(x)=(1/2)x^{T}Qx,$
where $Q\in{\mathbf{S}}_{+}^{d\times d}$ is a positive semidefinite matrix whose eigenvalues are contained in $\mu,L]$. For scalability reasons, we solve (DRO-PEP) with the function class $\mathcal{F}_{\mu,L}$ of $L$-smooth $\mu$-strongly convex functions, while we make sure to compute the worst-case performance in PEPit [8] with the tightest function class of convex quadratic functions $\mathcal{Q}_{\mu,L}$.
Algorithms
We compare gradient descent with step size $\eta>0$,
$x^{k+1}=x^{k}-\eta\nabla f(x^{k}),\qquad k=0,1,\dots,$ (GD)
and Nesterov’s fast gradient method [48][17, Section B.1.3],
$\begin{array}[t]{ll}x^{k+1}&=y^{k}-\eta\nabla f(y^{k})\\ y^{k+1}&=x^{k+1}+\beta_{k}(x^{k+1}-x^{k}),\qquad k=0,1,\dots,\end{array}$ (FGM)
where $\eta\in(0,1/L]$. The initial iterate $x^{0}\in{\mathbf{R}}^{d}$ satisfies the condition $\|x^{0}-x^{\star}\|\leq r$}, and $y^{0}=x^{0}$. Unless otherwise stated, we recursively define the momentum coefficient $\beta_{k}$ as [17, Section B.1.3]
$\begin{array}[t]{ll}\alpha_{k+2}&\leftarrow\;\text{Solve }(\alpha_{k+2}-\alpha_{k+1})^{2}-\alpha_{k+2}-q\alpha_{k+2}^{2}=0\\ \beta_{k}&\,=\;\dfrac{(\alpha_{k+2}-\alpha_{k+1})(\alpha_{k+1}(1-q)-\alpha_{k}-1)}{\alpha_{k+2}(2q\alpha_{k+1}+1)-q\alpha^{2}_{k+1}},\qquad k=0,1,\dots,\end{array}$
where $q=\mu/L$, $\alpha_{0}=0$, and $\alpha_{1}=(1-q)^{-1}$. Note that the quadratic equation for $\alpha_{k+2}\geq 0$ always has a closed-form solution.
Problem distribution
We sample values of $Q$, whose eigenvalues are all in $\mu,L]\subseteq{\mathbf{R}}$, to match existing results on average-case analysis [63, 16, 53]. Let $n=(\sqrt{L}-\sqrt{\mu})^{2}/(\sqrt{L}+\sqrt{\mu})^{2}$ and $\sigma=(\sqrt{L}+\sqrt{\mu})/2$. Define $A\in{\mathbf{R}}^{d\times nd}$ whose entries are i.i.d. Gaussian with zero mean and variance $\sigma^{2}$. In the regime of $d\to\infty$, the average value of the eigenvalues of $Q=(1/d)A^{T}A\in{\mathbf{R}}^{nd\times nd}$ follows the Marchenko-Pastur (MP) distribution [45] $\mu_{\text{MP}}$ whose probability density function is
$d\mu_{\text{MP}}(\lambda)=\frac{\sqrt{(L-\lambda)(\lambda-\mu)}}{2\pi n\sigma^{2}\lambda}1_{\lambda\in[\mu,L]}d\lambda,$
with $1_{[\mu,L]}$ being the 0-1 indicator function of the interval $\mu,L]$. In practice, we use sufficiently large $d$ and rejection sampling to ensure all sample values of $Q$ have eigenvalues in our predefined range $\mu,L]$.
Problem setup
Consider the initial condition $\|x^{0}-x^{\star}\|\leq r$. We set $d=300$, $L=25$, $r=10$, $\eta=1/L$, and sample $x^{0}$ from the sphere of radius $r$ centered at $x^{\star}=0$. We consider two alternative cases for $\mu$: $\mu=1$ for a strongly convex example and $\mu=0$ for a non-strongly convex example. In both cases, we used the distance-to-solution $\|x^{K}-x^{\star}\|^{2}$ as the performance metric. Later, we also consider the function-value gap $f(x^{K})-f^{\star}$ to compare our bounds to known results [16, 53] for this specific distribution.
Results
We present the plots showing convergence bounds for $\phi^{K}(f,x^{0})=\|x^{K}-x^{\star}\|^{2}$ vs. iteration number $K$ in Figure 3 for the strongly convex case ($\mu>0$) and Figure 4 for the non-strongly convex case ($\mu=0$). We summarize the fitted rates in Table 5 for the strongly convex case and Table 6 for the non-strongly convex case. In the case with $\mu=0$, we observe that the worst-case bound from PEP implies that neither GD nor FGM converge in the distance-to-solution metric. However, our DRO formulations show with high confidence that the iterates of both algorithms do converge. We further emphasize that the $\mathbf{CVaR}$ convergence bound with $\alpha=0.1$ is much closer to the average-case convergence bound than the worst-case bound, meaning that the true performances of both GD and FGM are close to their expected behavior for at least $(1-\alpha)$-fraction of problem instances. The timing results are available in Tables 10 and  11 in Appendix D.
Figure 3: Results for the unconstrained quadratic minimization experiment with strong convexity parameter $\mu=1$ and smoothness parameter $L=25$.
Table 5: Convergence rates of the unconstrained quadratic minimization experiment with $\mu=1$ and $L=25$. We generate samples corresponding to the strongly convex quadratic minimization problem of [16].
$O(\rho^{K}K^{-\gamma})$-convergence rates GD FGM
$\rho$ $\gamma$ $\rho$ $\gamma$
theoretical worst-case $1-(\mu/L)$ 0.0 $1-\sqrt{\mu/L}$ 0.0
theoretical worst-case value 0.96 0.0 0.8 0.0
fitted $\mathbf{CVaR}_{\alpha}(\|x^{K}-x^{\star}\|^{2})$ value 0.9155 0.7173 0.6870 0.0
fitted $\mathop{\bf E{}}(\|x^{K}-x^{\star}\|^{2})$ value 0.9060 0.8440 0.6709 0.2051
Figure 4: Results for the quadratic minimization experiment with strong convexity parameter $\mu=0$ and smoothness parameter $L=25$. The worst-case bounds for both algorithms remain constant since the algorithms do not converge in terms of the distance to optimality metric for this function class.
Table 6: Convergence rates in terms of $\|x^{K}-x^{\star}\|^{2}$ of the unconstrained quadratic minimization experiment with $\mu=0$ and $L=25$. We generate samples corresponding to the non-strongly convex quadratic minimization problem of [16].
$O(\rho^{K}K^{-\gamma})$-convergence rates GD FGM
$\rho$ $\gamma$ $\rho$ $\gamma$
theoretical worst-case 1.0 0.0 1.0 0.0
fitted $\mathbf{CVaR}_{\alpha}(\|x^{K}-x^{\star}\|^{2})$ value 1.0 0.5275 1.0 0.7820
fitted $\mathop{\bf E{}}(\|x^{K}-x^{\star}\|^{2})$ value 1.0 0.5422 0.9919 0.7553
Comparison with known average-case rates
As discussed in Section 2.1, probabilistic analysis can be challenging as it requires full knowledge of the underlying distribution $\mathbf{P}$}. One such analysis derives the average-case convergence rate in terms of function-value gap $\mathop{\bf E{}}_{(f,x^{0})\sim{\mathbf{P}}}(f(x^{K})-f^{\star})$, when the non-strongly convex quadratic objective function follows the MP distribution and the initial iterate satisfies $\mathop{\bf E{}}_{x^{0}}\big((x^{0}-x^{\star})(x^{0}-x^{\star})^{T}\big)=r^{2}I$ [16, 53]. In such cases, the average-case convergence rate of GD is of order $O(K^{-1.5})$ and that of FGM with momentum coefficient $\beta_{k}=k/(k+3)$ is of order $O(K^{-3}\log K)$.
We numerically compare our fitted (DRO-PEP) rates in terms of function-value gap with known rates [16, 53]. For our curve-fitting procedure, we fix $\rho=1$ and consider an extra $\log(K+1)$-term to the fit in Section 3.6 to match the logarithmic term in the average-case rate of FGM. That is, we consider the following fit:
$\log\phi_{K}\approx-\gamma\log(K+1)+\omega\log\big(\log(K+1)\big)+\log C,$
where $\gamma\geq 0$, $\omega\geq 0$, and $C>0$ are least-squares problem variables.
We set the same parameters as Figures 3 and 4, except for $\mu=0$ and $L=1$. We present the results in Figure 5 with a rate summary in Table 7. For both GD and FGM, we find that the fitted rate constants are close to the theoretical values. This experiment shows our method is able to approximately recover known theoretical rates without any knowledge of the underlying MP distribution. The timing results are available in Table 12 in Appendix D.
Figure 5: Results for the quadratic minimization experiment with strong convexity parameter $\mu=0$ and smoothness parameter $L=1$ in terms of objective value. Fitting the known form of the average rate when the quadratic terms are sampled from the MP distribution to the numerical values from (DRO-PEP) show a matching fit.
Table 7: Convergence rates of the unconstrained quadratic minimization experiment with $\mu=0$ and $L=1$ following the MP distribution (theoretical analysis in [16, 53]). For both fits, we enforce $\rho=1$.
$O(K^{-\gamma}(\log K)^{\omega})$-convergence rates GD FGM
$\gamma$ $\omega$ $\gamma$ $\omega$
theoretical worst-case $1.0$ $0.0$ $2.0$ $0.0$
theoretical $\mathop{\bf E{}}(f(x^{K})-f^{\star})$ value $1.5$ $0.0$ $3.0$ $1.0$
fitted $\mathop{\bf E{}}(f(x^{K})-f^{\star})$ value $1.478$ $0.0$ $2.819$ $0.9134$
4.2 Logistic regression with regularization
Consider a binary classification problem, with given data $\{(a_{i},b_{i})\}_{i=1}^{n}$ where $a_{i}\in{\mathbf{R}}^{d}$ are the feature vectors and $b_{i}\in\{0,1\}$ are boolean classes. The goal is to learn a feature vector $x\in{\mathbf{R}}^{d}$ from training data that linearly models the log-odds of a feature belonging to one of the classes. For this experiment, we add an $\ell_{2}$ regularizer with parameter $\lambda>0$. The associated objective function becomes
$f(x)=-\frac{1}{n}\left(\sum_{i=1}^{n}b_{i}\log\sigma(a_{i}^{T}x)+(1-b_{i})\log(1-\sigma(a_{i}^{T}x))\right)+\frac{\lambda}{2}\|x\|_{2}^{2},$
where $\sigma(z)=1/(1+e^{-z})$ is the sigmoid function [9, Chapter 7].
Algorithms
We again compare gradient descent and Nesterov’s fast gradient method as described in Section 4.1.
Problem distribution
Let $A\in{\mathbf{R}}^{n\times d}$ be the matrix of data points $\{a_{i}\}_{i=1}^{n}\subseteq{\mathbf{R}}^{d}$ stacked as row vectors and $b\in{\mathbf{R}}^{n}$ be the stacked vector of classes. Also, let ${\lambda_{\rm max}}(A^{T}A)$ be the maximum eigenvalue of $A^{T}A$. From [4, Theorem 5.12], it can be shown that the smoothness parameter is $L={\lambda_{\rm max}}(A^{T}A)/(4n)+\lambda$ and the strong convexity parameter is $\mu=\lambda$.
To sample a problem instance, we generate elements of $A$ i.i.d. from the Gaussian distribution $\mathcal{N}(0,\sigma_{A}^{2})$ for all except the last column, which we set to be all 1s. We also sample a feature vector $\tilde{x}\in{\mathbf{R}}^{d}$ with $\tilde{x}_{i}$ sampled from the uniform distribution $\mathcal{U}[-\tilde{x}^{\rm max},\tilde{x}^{\rm max}]$ with probability $p$ and 0 otherwise. To generate $b$, we sample a noise term $\varepsilon\in{\mathbf{R}}^{n}$ from a standard Gaussian and set $b_{i}=1$ if $\sigma(a_{i}^{T}\tilde{x}+\varepsilon_{i})>0$ and 0 otherwise. While sampling problems for our experiment, we tracked the largest value of $L$ across all sample problems and use this for the PEP comparison.
Problem setup
Consider the initial condition $\|x^{0}-x^{\star}\|\leq r$. We set $n=1000,~d=50,~p=0.3,~\sigma_{A}=4,~r=1,~\tilde{x}^{\rm max}=3,~\eta=1/L,~\lambda=10^{-2}$, and sample $x^{0}$ from the sphere of radius $r$ centered at $x^{\star}=0$. For our experiments, we found $L=6.300$ and chose function-value gap $f(x^{K})-f^{\star}$ as the performance metric.
Results
The plots showing convergence bounds for $\phi^{K}(f,x^{0})=f(x^{K})-f^{\star}$ vs. iteration number $K$ are presented in Figure 6 and fitted curve rates are available in Table 8. As in Section 4.1, our (DRO-PEP) results are able to much more closely match the out-of-sample behavior with an order of magnitude improvement over the worst-case by $K=30$. The timing results are available in Table 13 in Appendix D.
Figure 6: Results for the logistic regression experiment with GD and FGM.
Table 8: Convergence rates of the logistic regression experiment for GD and FGM, in terms of the worst-case, $\mathbf{CVaR}_{\alpha}$ with $\alpha=0.1$, and the average-case $f(x^{K})-f^{\star}$.
$O(\rho^{K}K^{-\gamma})$-convergence rates GD FGM
$\rho$ $\gamma$ $\rho$ $\gamma$
theoretical worst-case $1-(\mu/L)$ 0.0 $1-\sqrt{\mu/L}$ 0.0
theoretical worst-case value 0.9984 0.0 0.9602 0.0
fitted $\mathbf{CVaR}_{\alpha}(\phi^{K}(f,x^{0}))$ value 0.9416 0.7223 0.9160 1.2447
fitted $\mathop{\bf E{}}(\phi^{K}(f,x^{0}))$ value 0.9185 0.5860 0.8892 1.2039
4.3 Lasso
Consider an $\ell_{1}$-regularized least squares problem, commonly referred to as the Lasso [72]. The objective of this problem is
$f(x)=(1/2)\|Ax-b\|_{2}^{2}+\lambda\|x\|_{1},$ (13)
with $A\in{\mathbf{R}}^{m\times d}$, $b\in{\mathbf{R}}^{m}$, and regularization parameter $\lambda>0$.
Algorithms
As the objective function (13) is non-differentiable, we analyze different proximal methods [54], including the Iterative Soft-thresholding Algorithm (ISTA), and the Fast Iterative Soft-thresholding Algorithm (FISTA) [5]. Given parameter $\delta>0$, the (scalar) soft-thresholding function is given by $\mathcal{S}_{\delta}(v)=\min\left\{v+\delta,\,\max\left\{v-\delta,\,0\right\}\right\}$, whose vector version is element-wise application of $\mathcal{S}_{\delta}$. The ISTA iterations with step size $\eta>0$ are defined by
$x^{k+1}=\mathcal{S}_{\lambda\eta}(x^{k}-\eta A^{T}(Ax^{k}-b)),\qquad k=0,1,\dots.$ (ISTA)
The FISTA iterations with step size $\eta>0$ are given by
$\begin{array}[]{lll}x^{k+1}&=\mathcal{S}_{\lambda\eta}(y^{k}-\eta A^{T}(Ay^{k}-b))\\ y^{k+1}&=x^{k+1}+\beta_{k}(x^{k+1}-x^{k}),&\qquad k=0,1,\dots,\end{array}$ (FISTA)
with initialization $x^{0}=y^{0}$ and momentum coefficient $\beta_{k}=(\alpha_{k}-1)/\alpha_{k+1}$, where $\{\alpha_{k}\}$ is a sequence of scalars recursively defined as $\alpha_{k+1}=(1+\sqrt{1+4\alpha_{k}^{2}})/2$ from $\alpha_{0}=1$.
Problem distribution
We use a sparse coding example, which is the problem of recovering a sparse vector $\tilde{x}\in{\mathbf{R}}^{d}$ from its noisy linear measurements $b=A\tilde{x}+\varepsilon\in{\mathbf{R}}^{n}$, where $A\in{\mathbf{R}}^{n\times d}$ is a known dictionary matrix and $\varepsilon\in{\mathbf{R}}^{n}$ with $\varepsilon_{i}\sim\mathcal{N}(0,\sigma_{\varepsilon}^{2})$ is the noise [14]. We sampled one value of $A$ for the entire experiment (both training and testing) and construct a distribution of instances by sampling values of $\tilde{x}$ and $\varepsilon$ to form $b$. Similarly to [14, Section 4.1], we sampled the entries of $A$ i.i.d. from the Gaussian distribution $\mathcal{N}(0,1/n)$ with $40\%$ nonzeros and normalized its columns to have unit norm. After generating $A$, we experimentally computed the values of the smoothness parameter $L$ and the initial distance radius $r$ to use for the PEP comparison.
Problem setup
[…]
Results
The plots showing convergence bounds for $\phi^{K}(f,x^{0})=f(x^{K})-f^{\star}$ vs. $K$ are presented in Figure 7 and fitted curve rates are available in Table 9. For this distribution, we observe that ISTA exhibits a generally sublinear convergence trend. Although ISTA displays contracting behavior, we suspect that fitted contraction factors being slightly less than $1$ is an artifact of applying the fitting scheme over a finite horizon of algorithm runs. In contrast, FISTA’s behavior more closely resembles linear convergence, and our fitted rates are able to capture this phenomenon. The timing results are available in Table 14 in Appendix D.
Figure 7: Results for the Lasso minimization experiment with ISTA and FISTA.
Table 9: Convergence rates of the Lasso minimization experiment for ISTA and FISTA, in terms of the worst-case, $\mathbf{CVaR}_{\alpha}$, and the average-case $f(x^{K})-f^{\star}$.
$O(\rho^{K}K^{-\gamma})$-convergence rates ISTA FISTA
$\rho$ $\gamma$ $\rho$ $\gamma$
theoretical worst-case 1.0 1.0 1.0 2.0
fitted $\mathbf{CVaR}_{\alpha}(f(x^{K})-f^{\star})$ value 0.9951 1.1326 0.8396 0.0
fitted $\mathop{\bf E{}}(f(x^{K})-f^{\star})$ value 0.9737 0.7224 0.8064 0.0
5 Conclusion
In this paper, we presented a probabilistic performance estimation framework to quantify the convergence behavior of deterministic first-order methods applied on a distribution of convex optimization problems. By combining performance estimation problem (PEP) and distributionally robust optimization (DRO) techniques, we derive data-driven probabilistic performance guarantees that only require iterate trajectories from a finite set of sampled function instances. By fitting the resulting bounds, we also construct data-driven convergence rates that consistently improve worst-case analysis. To the best of our knowledge, this is the first work to apply PEP to establish probabilistic convergence rates. A key limitation of our approach is the scalability of (DRO-PEP), whose dimension is proportional to the number of samples and the square of the number of iterations. One way to address this issue, is to adopt scenario reduction techniques, such as mean robust optimization [76], that cluster the data points together, effectively reducing the size of the DRO problem.
Appendix A Proof of Theorem 1
The following theorem states that every $\mathcal{F}\subseteq\mathcal{F}_{0,\infty}$ in Table 2 is a closed subspace of the Sobolev space $W_{loc}^{\infty}$ with Sobolev norm (5), when restricted to the compact domain $\mathcal{D}\subseteq{\mathbf{R}}^{d}$.
Theorem 4.
Consider a class of closed, lower-semicontinuous, and convex functions $\mathcal{F}_{0,\infty}$ with compact domain $\mathcal{D}\subseteq{\mathbf{R}}^{d}$. Then, $\mathcal{F}_{0,\infty}$ is a closed subset of Sobolev space $W_{loc}^{\infty}$.
Proof.
Note that every $f\in\mathcal{F}_{0,\infty}$ is locally Lipschitz, which further implies that $\nabla f$ is bounded on compact domain $\mathcal{D}$. Therefore, the Sobolev norm (5) of $f$ on $\mathcal{D}$ is
$\big\|f\big\|=\sup_{x\in\mathcal{D}}\big|f(x)\big|+\sup_{x\in\mathcal{D}}\big\|\nabla f(x)\big\|<\infty,$
which implies that $\mathcal{F}_{0,\infty}\subseteq W_{loc}^{\infty}$. Since $\mathcal{F}_{0,\infty}$ is a convex cone, it is a closed subset of Sobolev space $W_{loc}^{\infty}$. ∎
Lemma 2.
Given closed, lower-semicontinuous, and convex functions $f_{1},f_{2}$ over a compact domain $\mathcal{D}\subseteq{\mathbf{R}}^{d}$, let $\varepsilon_{f}=\sup_{x\in\mathcal{D}}|f_{1}(x)-f_{2}(x)|$ and $\varepsilon_{g}=\sup_{x\in\mathcal{D}}\big\|\nabla f_{1}(x)-\nabla f_{2}(x)\big\|$. For any $x_{1},x_{2}\in\mathcal{D}$ with $\varepsilon_{x}=\|x_{1}-x_{2}\|$, we get
$\big|f_{1}(x_{1})-f_{2}(x_{2})\big|$ $\leq\left(\sup_{(f,x)\in\mathcal{F}\times\mathcal{D}}\,\big\|\nabla f(x)\big\|\right)\varepsilon_{x}+\varepsilon_{f}$
$\big\|\nabla f_{1}(x_{1})-\nabla f_{2}(x_{2})\big\|$ $\leq\left(\sup_{(f,x)\in\mathcal{F}\times\mathcal{D}}\,\textnormal{Lip}(\nabla f)\right)\varepsilon_{x}+\varepsilon_{g}.$
Proof.
We have
$\big|f_{1}(x_{1})-f_{2}(x_{2})\big|$ $\leq\big|f_{1}(x_{1})-f_{1}(x_{2})\big|+\big|f_{1}(x_{2})-f_{2}(x_{2})\big|$
$\leq\left(\sup_{x\in\mathcal{D}}\big\|\nabla f_{1}(x)\big\|\right)\big\|x_{1}-x_{2}\big\|+\big|f_{1}(x_{2})-f_{2}(x_{2})\big|$
$\leq\left(\sup_{(f,x)\in\mathcal{F}\times\mathcal{D}}\,\big\|\nabla f(x)\big\|\right)\varepsilon_{x}+\varepsilon_{f},$
and
$\big\|\nabla f_{1}(x_{1})-\nabla f_{2}(x_{2})\big\|$ $\leq\big\|\nabla f_{1}(x_{1})-\nabla f_{1}(x_{2})\big\|+\big\|\nabla f_{1}(x_{2})-\nabla f_{2}(x_{2})\big\|$
$\leq\Big(\text{Lip}(\nabla f_{1})\Big)\big\|x_{1}-x_{2}\big\|+\big\|\nabla f_{1}(x_{2})-\nabla f_{2}(x_{2})\big\|$
$\leq\left(\underset{(f,x)\in\mathcal{F}\times\mathcal{D}}{\sup}\,\textnormal{Lip}(\nabla f)\right)\varepsilon_{x}+\varepsilon_{g}.$
We now prove Theorem 1.
Proof of Theorem 1.
From Lemma 2, the definition of the Sobolev norm $\|\cdot\|$ immediately gives $\|f_{1}-f_{2}\|=\varepsilon_{g}+\varepsilon_{f}$ and $\|(f_{1},x_{1}^{0})-(f_{2},x_{2}^{0})\|=\varepsilon_{x}+\varepsilon_{g}+\varepsilon_{f}$. Define $P_{i}\in{\mathbf{R}}^{d\times(K+2)}$ as in (4), which gives $G_{i}=P_{i}^{T}P_{i}$. Then
$\big\|Z_{1}-Z_{2}\big\|^{2}$ $=\big\|G_{1}-G_{2}\big\|_{F}^{2}+\big\|F_{1}-F_{2}\big\|^{2}$
$=\big\|P_{1}^{T}P_{1}-P_{2}^{T}P_{2}\big\|_{F}^{2}+\big\|F_{1}-F_{2}\big\|^{2}$
$\leq\big\|P_{1}+P_{2}\big\|_{\text{op}}^{2}\big\|P_{1}-P_{2}\big\|_{F}^{2}+\big\|F_{1}-F_{2}\big\|^{2},$
with $\|\cdot\|_{\text{op}}$ being the operator norm. For notational simplicity, let $L=\sup_{(f,x)\in\mathcal{F}}\,\big\|\nabla f(x)\big\|$.
First of all,
$\big\|P_{1}+P_{2}\big\|_{\text{op}}\leq 2\max\left\{\text{diam}(\mathcal{D}),\;\sup_{(f,x)\in\mathcal{H}}\big\|\nabla f(x)\big\|\right\}=2\,\text{diam}(\mathcal{D})\,\max\{1,L\}=C,$
holds true for all $(f_{1},x_{1}^{0}),(f_{2},x_{2}^{0})\in\mathcal{H}$. Note that given a convergent algorithm $\mathcal{A}$, $\big\|x_{1}^{k}-x_{2}^{k}\big\|\leq\big\|x_{1}^{0}-x_{2}^{0}\big\|$ holds for any $k$. Furthermore, from
$\big\|P_{1}-P_{2}\big\|_{F}^{2}=\big\|x_{1}^{0}-x_{2}^{0}\big\|^{2}+{\textstyle\sum_{k=0}^{K}}\big\|g_{1}^{k}-g_{2}^{k}\big\|^{2},$
and Lemma 2, we get
$\big\|(G_{1},F_{1})-(G_{2},F_{2})\big\|^{2}$
$\leq C^{2}\,\big\|x_{1}^{0}-x_{2}^{0}\big\|^{2}+C^{2}\,{\textstyle\sum_{k=0}^{K}}\big\|g_{1}^{k}-g_{2}^{k}\big\|^{2}+{\textstyle\sum_{k=0}^{K}}(f_{1}^{k}-f_{2}^{k})^{2}$
$\leq C^{2}\varepsilon_{x}^{2}+C^{2}{\textstyle\sum_{k=0}^{K}}\left(L\varepsilon_{x}+\varepsilon_{g}\right)^{2}+{\textstyle\sum_{k=0}^{K}}\left(L\,\mathrm{diam}(\mathcal{D})\,\varepsilon_{x}+\varepsilon_{f}\right)^{2}$
$\leq\left\{C^{2}+C^{2}(K+1)^{2}\max\{1,L\}^{2}+\max\{1,L\}^{2}(K+1)^{2}\,\mathrm{diam}\,(\mathcal{D})^{2}\right\}^{2}(\varepsilon_{x}+\varepsilon_{g}+\varepsilon_{f})^{2}$
$=C_{K}^{2}\big\|(f_{1},x_{1}^{0})-(f_{2},x_{2}^{0})\big\|^{2}.$
Appendix B Proof of Proposition 1
Proof of Proposition 1.
As we will see in Theorem 2, these worst-case expectations are equivalent to robust optimization problems, in the form of maximization problem. The constraint on the ambiguity set with radius $\varepsilon$ becomes the uncertainty set constraint
$(1/N){\textstyle\sum_{i=1}^{N}}\|Z_{i}-\widehat{Z}_{i}\|\leq\varepsilon,$
in terms of optimization variables $Z_{i}\in\mathcal{F}\cap\mathcal{X}$ for $i=1,\dots,N$. Similarly, the ambiguity set $\mathcal{P}$ over the distributions supported on $\mathcal{H}$ equivalently writes as the uncertainty set constraint
$(1/N){\textstyle\sum_{i=1}^{N}}\|(f_{i},x_{i}^{0})-(\hat{f}_{i},\hat{x}_{i}^{0})\|\leq\varepsilon,$
in terms of variables $(f_{i},x_{i}^{0})\in\mathcal{H}$ for $i=1,\dots,N$.
According to Theorem 1, $(1/N){\textstyle\sum_{i=1}^{N}}\|(f_{i},x_{i}^{0})-(\hat{f}_{i},\hat{x}_{i}^{0})\|\leq(C_{K})^{-1}\varepsilon$ implies
$(1/N){\textstyle\sum_{i=1}^{N}}\|Z_{i}-\widehat{Z}_{i}\|\leq(1/N){\textstyle\sum_{i=1}^{N}}C_{K}\|(f_{i},x_{i}^{0})-(\hat{f}_{i},\hat{x}_{i}^{0})\|\leq\varepsilon.$
As the worst-case expectation in terms of $Z=(G,F)$ is a maximization problem over larger feasible set (or ambiguity set), we get the desired inequality between worst-case expectations. ∎
Appendix C Proofs of Theorem 2 and Corollary 2.1
We derive a convex formulation of the right-hand size of (8) by adopting the proof strategies from [46, Theorem 4.2] and [76, Appendix A]. We first rewrite it to highlight the support constraints
$\sup_{\mathbf{Q}_{Z}\in B_{{\varepsilon}}}\,\inf_{\ell\in\mathcal{L}}\,\mathop{\bf E{}}_{Z\sim\mathbf{Q}_{Z}}\,\left(\ell(Z)\right)\;=\;\begin{array}[t]{ll}\sup&\inf_{\ell\in\mathcal{L}}\,\mathop{\bf E{}}_{(G,F)\sim\mathbf{Q}_{Z}}\,\left(\ell(G,F)\right)\\ \text{s.t.}&\mathrm{supp}\,\mathbf{Q}_{Z}=\mathcal{F}\cap\mathcal{X}\\ &\mathbf{Q}_{Z}\in B_{{\varepsilon}}.\end{array}$ (14)
We show that this problem is equivalent to the following minimization problem.
Lemma 3.
Consider the setup of Theorem 2. For every $\varepsilon>0$,
$\begin{array}[t]{lll}\underset{\mathbf{Q}_{Z}\in B_{{\varepsilon}}}{\sup}&\underset{\ell\in\mathcal{L}}{\inf}&\underset{(G,F)\sim\mathbf{Q}_{Z}}{\mathop{\bf E{}}}\,\left(\ell(G,F)\right),\end{array}$
$=\begin{array}[t]{lll}\textnormal{minimize}&(1/N)\sum_{i=1}^{N}s_{i}\\ \textnormal{subject to}&c_{\mathrm{obj}}^{j}t-c_{0}\tau_{i}^{j}-\big\langle(X_{i}^{j},Y_{i}^{j}),(\widehat{G}_{i},\widehat{F}_{i})\big\rangle+\lambda\varepsilon\leq s_{i}\\ &-\mathcal{S}^{*}(y_{i}^{j})-(X_{i}^{j},Y_{i}^{j})+\tau_{i}^{j}(A_{0},b_{0})-(A_{\mathrm{obj}}^{j},b_{\mathrm{obj}}^{j})\in{\mathbf{S}}_{+}^{K+2}\times\{0\}\\ &\big\|(X_{i}^{j},Y_{i}^{j})\big\|_{*}\leq\lambda,\hfill i=1,\dots,N,\;j=1,\dots,J,\end{array}$
with variables $s_{i}\in{\mathbf{R}}$, $t\in{\mathbf{R}}$, $\tau_{i}^{j}\in{\mathbf{R}}_{+}$, $y_{i}^{j}\in{\mathbf{R}}_{+}^{|M|}$, $(X_{i}^{j},Y_{i}^{j})\in{\mathbf{S}}^{K+2}\times{\mathbf{R}}^{K+1}$, and $\lambda\in{\mathbf{R}}_{+}$.
Proof.
We follow the proof structure in [46, Theorem 4.2] to derive the finite-dimensional dual problem to the worst-case expectation in (8), which is an infinite-dimensional optimization problem.
$\begin{array}[t]{ll}\underset{\mathbf{Q}_{Z}\in B_{{\varepsilon}}}{\sup}&\underset{\ell\in\mathcal{L}}{\inf}\;\underset{(G,F)\sim\mathbf{Q}_{Z}}{\mathop{\bf E{}}}\,\left(\ell(G,F)\right)\end{array}$
$=\begin{array}[t]{lll}\underset{\mathbf{Q}_{Z}^{i}\in\mathcal{M}(\mathcal{F}\cap\mathcal{X})}{\textnormal{maximize}}&\underset{\lambda\in{\mathbf{R}}_{+}}{\textnormal{minimize}}&\underset{\ell\in\mathcal{L}}{\inf}\;\Bigg(\underset{(G,F)\sim\mathbf{Q}_{Z}}{\mathop{\bf E{}}}\;\left(\ell(G,F)\right)\\ &&\qquad\quad-\lambda\left((1/N)\sum_{i=1}^{N}\underset{(G,F)\sim\mathbf{Q}_{Z}^{i}}{\mathop{\bf E{}}}\,\left(\big\|(G,F)-(\widehat{G}_{i},\widehat{F}_{i})\big\|\right)-\varepsilon\right)\Bigg)\end{array}$
$=\begin{array}[t]{lll}\underset{\mathbf{Q}_{Z}^{i}\in\mathcal{M}(\mathcal{F}\cap\mathcal{X})}{\textnormal{maximize}}&\underset{\lambda\in{\mathbf{R}}_{+},\,\ell\in\mathcal{L}}{\textnormal{minimize}}&(1/N)\sum_{i=1}^{N}\underset{(G,F)\sim\mathbf{Q}_{Z}^{i}}{\mathop{\bf E{}}}\;\left(\ell(G,F)-\lambda\big\|(G,F)-(\widehat{G}_{i},\widehat{F}_{i})\big\|+\lambda\varepsilon\right)\end{array}$
$=\begin{array}[t]{lll}\underset{\mathbf{Q}_{Z}^{i}\in\mathcal{M}(\mathcal{F}\cap\mathcal{X})}{\textnormal{maximize}}&\underset{\lambda\in{\mathbf{R}}_{+},\,t\in{\mathbf{R}}}{\textnormal{minimize}}&(1/N)\sum_{i=1}^{N}\underset{(G,F)\sim\mathbf{Q}_{Z}^{i}}{\mathop{\bf E{}}}\;\left(\ell(G,F)-\lambda\big\|(G,F)-(\widehat{G}_{i},\widehat{F}_{i})\big\|+\lambda\varepsilon\right)\end{array}$
$\leq\begin{array}[t]{lll}\underset{\lambda\in{\mathbf{R}}_{+},\,t\in{\mathbf{R}}}{\textnormal{minimize}}&\underset{\mathbf{Q}_{Z}^{i}\in\mathcal{M}(\mathcal{F}\cap\mathcal{X})}{\textnormal{maximize}}&(1/N)\sum_{i=1}^{N}\underset{(G,F)\sim\mathbf{Q}_{Z}^{i}}{\mathop{\bf E{}}}\;\left(\ell(G,F)-\lambda\big\|(G,F)-(\widehat{G}_{i},\widehat{F}_{i})\big\|+\lambda\varepsilon\right)\end{array}$
$=\begin{array}[t]{ll}\underset{\lambda\in{\mathbf{R}}_{+},\,t\in{\mathbf{R}}}{\textnormal{minimize}}&(1/N)\sum_{i=1}^{N}\underset{(G_{i},F_{i})\in\mathcal{F}\cap\mathcal{X}}{\textnormal{maximize}}\;\left(\ell(G_{i},F_{i})-\lambda\big\|(G_{i},F_{i})-(\widehat{G}_{i},\widehat{F}_{i})\big\|+\lambda\varepsilon\right).\end{array}$
The first equality comes from rewriting the worst-case expectation problem. The second equality comes from the law of total distribution. The third equality comes from the parametrization of $\ell\in\mathcal{L}$ with parameter $t\in{\mathbf{R}}$. The following inequality above comes from the max-min inequality, and it is actually an equality for any $\varepsilon>0$, due to the strong duality of moment problems. The last equality comes from $\mathcal{M}(\mathcal{F}\cap\mathcal{X})$ containing all the Dirac distributions supported on $\mathcal{F}\cap\mathcal{X}$.
Dualizing the inner maximization problem, we get
$\begin{array}[t]{ll}\textnormal{minimize}&(1/N)\sum_{i=1}^{N}s_{i}\\ \textnormal{subject to}&[-\ell^{j}]^{*}(\mu_{i}^{j}-\nu_{i}^{j})+\sigma_{\mathcal{F}\cap\mathcal{X}}(\nu_{i}^{j})-\big\langle\mu_{i}^{j},(\widehat{G}_{i},\widehat{F}_{i})\big\rangle+\lambda\varepsilon\leq s_{i}\\ &\big\|\mu_{i}^{j}\big\|_{*}\leq\lambda,\hfill i=1,\dots,N,\,j=1,\dots,J,\end{array}$
with variables $s_{i}\in{\mathbf{R}}$, $\lambda\in{\mathbf{R}}_{+}$, $t\in{\mathbf{R}}$, $\mu_{i}^{j}\in{\mathbf{S}}^{K+2}\times{\mathbf{R}}^{K+1}$, and $\nu_{i}^{j}\in{\mathbf{S}}^{K+2}\times{\mathbf{R}}^{K+1}$. Note that $\ell^{j}]^{*}$ is a convex conjugate with respect to variables in ${\mathbf{S}}^{K+2}\times{\mathbf{R}}^{K+1}$, i.e.,
$[-\ell^{j}]^{*}(\mu_{i}^{j}-\nu_{i}^{j})$ $=\underset{(G,F)\in{\mathbf{S}}^{K+2}\times{\mathbf{R}}^{K+1}}{\sup}\;\left(\big\langle\mu_{i}^{j}-\nu_{i}^{j},(G,F)\big\rangle+\big\langle(A_{\mathrm{obj}}^{j},b_{\mathrm{obj}}^{j}),(G,F)\big\rangle+c_{\mathrm{obj}}^{j}t\right)$
$=\begin{cases}c_{\mathrm{obj}}^{j}t&\textnormal{if }\mu_{i}^{j}-\nu_{i}^{j}+(A_{\mathrm{obj}}^{j},b_{\mathrm{obj}}^{j})=0,\\ \infty&\textnormal{o.w}.\end{cases}$
The support function $\sigma_{\mathcal{F}\cap\mathcal{X}}$ of
[…]
writes as
$\sigma_{\mathcal{F}\cap\mathcal{X}}(\nu_{i}^{j})$ $=\underset{(G,F)\in\mathcal{F}\cap\mathcal{X}}{\sup}\;\big\langle\nu_{i}^{j},(G,F)\big\rangle$
$=\begin{array}[t]{ll}\textnormal{maximize}&\big\langle\nu_{i}^{j},(G,F)\big\rangle\\ \textnormal{subject to}&\mathcal{S}(G,F)\in{\mathbf{R}}_{+}^{|M|}\\ &-\big\langle(A_{0},b_{0}),(G,F)\big\rangle-c_{0}\in{\mathbf{R}}_{+}\end{array}$
$=\begin{array}[t]{ll}\textnormal{minimize}&-c_{0}\tau_{i}^{j}\\ \textnormal{subject to}&-\mathcal{S}^{*}(y_{i}^{j})+\tau_{i}^{j}(A_{0},b_{0})-\nu_{i}^{j}\in{\mathbf{S}}_{+}^{K+2}\times\{0\},\end{array}$
with variables $(G,F)\in{\mathbf{S}}_{+}^{K+2}\times{\mathbf{R}}^{K+1}$, $\tau_{i}^{j}\in{\mathbf{R}}_{+}$ and $y_{i}^{j}\in{\mathbf{R}}_{+}^{|M|}$. The last equality comes from the strong duality argument via Slater’s condition, which is valid for function classes $\mathcal{F}$ in Table 2 [70, Theorem 6]. The resulting minimization problem becomes
$\begin{array}[t]{lll}\textnormal{minimize}&(1/N)\sum_{i=1}^{N}s_{i}\\ \textnormal{subject to}&c_{\mathrm{obj}}^{j}t-c_{0}\tau_{i}^{j}-\big\langle(X_{i}^{j},Y_{i}^{j}),\,(\widehat{G}_{i},\widehat{F}_{i})\big\rangle+\lambda\varepsilon\leq s_{i}\\ &-\mathcal{S}^{*}(y_{i}^{j})-(X_{i}^{j},Y_{i}^{j})+\tau_{i}^{j}(A_{0},b_{0})-(A_{\mathrm{obj}}^{j},b_{\mathrm{obj}}^{j})\in{\mathbf{S}}_{+}^{K+2}\times\{0\}\\ &\big\|(X_{i}^{j},Y_{i}^{j})\big\|_{*}\leq\lambda,\hfill i=1,\dots,N,\,j=1,\dots,J,\end{array}$
with variables $s_{i}\in{\mathbf{R}}$, $\lambda\in{\mathbf{R}}_{+}$, $t\in{\mathbf{R}}$, $y_{i}^{j}\in{\mathbf{R}}_{+}^{|M|}$, $\tau_{i}^{j}\in{\mathbf{R}}_{+}$, and $(X_{i}^{j},Y_{i}^{j})\in{\mathbf{S}}^{K+2}\times{\mathbf{R}}^{K+1}$. ∎
Next, we take the dual of the minimization problem to reach the maximization problem.
Lemma 4.
Consider the setup of Theorem 2. For every $\varepsilon>0$, the weak duality result holds true for the following pair of primal and dual problems:
$\begin{array}[t]{lll}\textnormal{minimize}&(1/N)\sum_{i=1}^{N}s_{i}\\ \textnormal{subject to}&c_{\mathrm{obj}}^{j}t-c_{0}\tau_{i}^{j}-\big\langle(X_{i}^{j},Y_{i}^{j}),\,(\widehat{G}_{i},\widehat{F}_{i})\big\rangle+\lambda\varepsilon\leq s_{i}\\ &-\mathcal{S}^{*}(y_{i}^{j})-(X_{i}^{j},Y_{i}^{j})+\tau_{i}^{j}(A_{0},b_{0})-(A_{\mathrm{obj}}^{j},b_{\mathrm{obj}}^{j})\in{\mathbf{S}}_{+}^{K+2}\times\{0\}\\ &\big\|(X_{i}^{j},Y_{i}^{j})\big\|_{*}\leq\lambda,\qquad(i=1,\dots,N,\;j=1,\dots,J),\end{array}$
$\geq\begin{array}[t]{ll}\textnormal{maximize}&\sum_{i=1}^{N}\sum_{j=1}^{J}\alpha_{i}^{j}\big\langle(A_{\mathrm{obj}}^{j},b_{\mathrm{obj}}^{j}),\,(G_{i}^{j},F_{i}^{j})\big\rangle\\ \textnormal{subject to}&\sum_{i=1}^{N}\sum_{j=1}^{J}\alpha_{i}^{j}\big\|(G_{i}^{j},F_{i}^{j})-(\widehat{G}_{i},\widehat{F}_{i})\big\|\leq\varepsilon,\end{array}$
with variables $(G_{i}^{j},F_{i}^{j})\in\mathcal{F}\cap\mathcal{X}$ and $(\alpha_{i}^{j})\in\Gamma$ such that
$\Gamma=\left\{(\alpha_{i}^{j})\in{\mathbf{R}}_{+}^{N\times J}\;\left|\;\begin{array}[]{ll}\sum_{j=1}^{J}\alpha_{i}^{j}=1/N,&i=1,\dots,N\\ \sum_{j=1}^{J}\left(\sum_{i=1}^{N}\alpha_{i}^{j}\right)c_{\mathrm{obj}}^{j}=0\end{array}\right.\right\}$.
In addition, the strong duality holds therefore the above inequality holds with equality.
Proof.
Consider the minimization problem in Theorem 2. Using the auxiliary variable $\lambda_{i}^{j}\in{\mathbf{R}}$ for $i=1,\dots,N$ and $j=1,\dots,J$, the problem writes as
$\begin{array}[t]{ll}\textnormal{minimize}&(1/N)\sum_{i=1}^{N}s_{i}\\ \textnormal{subject to}&c_{\mathrm{obj}}^{j}t-c_{0}\tau_{i}^{j}-\big\langle(X_{i}^{j},Y_{i}^{j})\,\,(\widehat{G}_{i},\widehat{F}_{i})\big\rangle+\lambda\varepsilon\leq s_{i}\\ &-\mathcal{S}^{*}(y_{i}^{j})-(X_{i}^{j},Y_{i}^{j})+\tau_{i}^{j}(A_{0},b_{0})-(A_{\mathrm{obj}}^{j},b_{\mathrm{obj}}^{j})\in{\mathbf{S}}_{+}^{K+2}\times\{0\}\\ &\lambda_{i}^{j}=\lambda,\hfill(i=1,\dots,N,\,j=1,\dots,J),\end{array}$
with variables $s_{i}\in{\mathbf{R}}$, $\lambda\in{\mathbf{R}}_{+}$, $t\in{\mathbf{R}}$, $y_{i}^{j}\in{\mathbf{R}}_{+}^{|M|}$, $\tau_{i}^{j}\in{\mathbf{R}}_{+}$, and $\big(\lambda_{i}^{j},(X_{i}^{j},Y_{i}^{j})\big)\in{\mathcal{K}_{\text{SOC}}}$. The Lagrangian of this problem with dual variables
$\alpha_{i}^{j}\in{\mathbf{R}}_{+},\quad(G_{i}^{j},F_{i}^{j})\in{\mathbf{S}}_{+}^{K+2}\times{\mathbf{R}}^{K+1},\quad\varepsilon_{i}^{j}\in{\mathbf{R}},$
writes as
$\mathbf{L}\left(s_{i},\lambda,t,y_{i}^{j},\big(\lambda_{i}^{j},(X_{i}^{j},Y_{i}^{j})\big),\;\alpha_{i}^{j},(G_{i}^{j},F_{i}^{j}),\varepsilon_{i}^{j}\right)$
$=\frac{1}{N}\sum_{i=1}^{N}s_{i}+\sum_{i=1}^{N}\sum_{j=1}^{J}\,\alpha_{i}^{j}\left(c_{\mathrm{obj}}^{j}t-c_{0}\tau_{i}^{j}-\big\langle(X_{i}^{j},Y_{i}^{j}),\,(\widehat{G}_{i},\widehat{F}_{i})\big\rangle+\lambda\varepsilon-s_{i}\right)$
$\quad+\sum_{i=1}^{N}\sum_{j=1}^{J}\,\big\langle\mathcal{S}^{*}(y_{i}^{j})+(X_{i}^{j},Y_{i}^{j})-\tau_{i}^{j}(A_{0},b_{0})+(A_{\mathrm{obj}}^{j},b_{\mathrm{obj}}^{j}),(G_{i}^{j},F_{i}^{j})\big\rangle+\sum_{i=1}^{N}\sum_{j=1}^{J}\,\varepsilon_{i}^{j}(\lambda_{i}^{j}-\lambda)$
$=\sum_{i=1}^{N}s_{i}\left(\frac{1}{N}-\sum_{j=1}^{J}\alpha_{i}^{j}\right)+\lambda\sum_{i=1}^{N}\sum_{j=1}^{J}\left(\alpha_{i}^{j}\varepsilon-\varepsilon_{i}^{j}\right)+t\sum_{j=1}^{J}\left(\sum_{i=1}^{N}\alpha_{i}^{j}\right)c_{\mathrm{obj}}^{j}$
$\quad+\sum_{i=1}^{N}\sum_{j=1}^{J}\,(y_{i}^{j})^{T}\left(\mathcal{S}(G_{i}^{j},F_{i}^{j})\right)-\sum_{i=1}^{N}\sum_{j=1}^{J}\,\tau_{i}^{j}\left(\big\langle(A_{0},b_{0}),(G_{i}^{j},F_{i}^{j})\big\rangle+\alpha_{i}^{j}c_{0}\right)$
$\quad+\sum_{i=1}^{N}\sum_{j=1}^{J}\,\left(\lambda_{i}^{j}\varepsilon_{i}^{j}+\big\langle(X_{i}^{j},Y_{i}^{j}),(G_{i}^{j},F_{i}^{j})-\alpha_{i}^{j}(\widehat{G}_{i},\widehat{F}_{i})\big\rangle\right)+\sum_{i=1}^{N}\sum_{j=1}^{J}\big\langle(A_{\mathrm{obj}}^{j},b_{\mathrm{obj}}^{j}),(G_{i}^{j},F_{i}^{j})\big\rangle.$
The corresponding dual problem is
$\begin{array}[t]{lll}\textnormal{maximize}&\sum_{i=1}^{N}\sum_{j=1}^{J}\big\langle(A_{\mathrm{obj}}^{j},b_{\mathrm{obj}}^{j}),\,(G_{i}^{j},F_{i}^{j})\big\rangle\\ \textnormal{subject to}&\sum_{j=1}^{J}\alpha_{i}^{j}=1/N,&i=1,\dots,N\\ &\left(\sum_{i=1}^{N}\sum_{j=1}^{J}\alpha_{i}^{j}\right)\varepsilon-\sum_{i=1}^{N}\sum_{j=1}^{J}\varepsilon_{i}^{j}\geq 0\\ &\sum_{j=1}^{J}\left(\sum_{i=1}^{N}\alpha_{i}^{j}\right)c_{\mathrm{obj}}^{j}=0\\ &\mathcal{S}(G_{i}^{j},F_{i}^{j})\in{\mathbf{R}}_{+}^{|M|},&i=1,\dots,N,\;j=1,\dots,J\\ &\big\langle(A_{0},b_{0}),\,(G_{i}^{j},F_{i}^{j})\big\rangle+\alpha_{i}^{j}c_{0}\leq 0,&i=1,\dots,N,\;j=1,\dots,J\\ &\left(\varepsilon_{i}^{j},\,(G_{i}^{j},F_{i}^{j})-\alpha_{i}^{j}(\widehat{G}_{i},\widehat{F}_{i})\right)\in{\mathcal{K}_{\text{SOC}}},&i=1,\dots,N,\;j=1,\dots,J,\end{array}$
$=\begin{array}[t]{lll}\textnormal{maximize}&\sum_{i=1}^{N}\sum_{j=1}^{J}\big\langle(A_{\mathrm{obj}}^{j},b_{\mathrm{obj}}^{j}),\,(G_{i}^{j},F_{i}^{j})\big\rangle\\ \textnormal{subject to}&\sum_{j=1}^{J}\alpha_{i}^{j}=1/N,&i=1,\dots,N\\ &\sum_{i=1}^{N}\sum_{j=1}^{J}\big\|(G_{i}^{j},F_{i}^{j})-\alpha_{i}^{j}(\widehat{G}_{i},\widehat{F}_{i})\big\|\leq\varepsilon\\ &\sum_{j=1}^{J}\left(\sum_{i=1}^{N}\alpha_{i}^{j}\right)c_{\mathrm{obj}}^{j}=0\\ &\mathcal{S}(G_{i}^{j},F_{i}^{j})\in{\mathbf{R}}_{+}^{|M|},&i=1,\dots,N,\;j=1,\dots,J,\\ &\big\langle(A_{0},b_{0}),\,(G_{i}^{j},F_{i}^{j})\big\rangle+\alpha_{i}^{j}c_{0}\leq 0,&i=1,\dots,N,\;j=1,\dots,J,\end{array}$
with variables $(G_{i}^{j},F_{i}^{j})\in{\mathbf{S}}_{+}^{K+2}\times{\mathbf{R}}^{K+1}$ and $\alpha_{i}^{j}\in{\mathbf{R}}_{+}$. Note that $\varepsilon_{i}^{j}$ is removed using $\big\|(G_{i}^{j},F_{i}^{j})-\alpha_{i}^{j}(\widehat{G}_{i},\widehat{F}_{i})\big\|\leq\varepsilon_{i}^{j}$. Using the change of variables $(G_{i}^{j},F_{i}^{j})\leftarrow\alpha_{i}^{j}(G_{i}^{j},F_{i}^{j})$, we get the problem
$\begin{array}[t]{lll}\textnormal{maximize}&\sum_{i=1}^{N}\sum_{j=1}^{J}\alpha_{i}^{j}\big\langle(A_{\mathrm{obj}}^{j},b_{\mathrm{obj}}^{j}),\,(G_{i}^{j},F_{i}^{j})\big\rangle\\ \textnormal{subject to}&\sum_{j=1}^{J}\alpha_{i}^{j}=1/N,&i=1,\dots,N\\ &\sum_{i=1}^{N}\sum_{j=1}^{J}\alpha_{i}^{j}\big\|(G_{i}^{j},F_{i}^{j})-(\widehat{G}_{i},\widehat{F}_{i})\big\|\leq\varepsilon\\ &\sum_{j=1}^{J}\left(\sum_{i=1}^{N}\alpha_{i}^{j}\right)c_{\mathrm{obj}}^{j}=0\\ &\alpha_{i}^{j}\mathcal{S}(G_{i}^{j},F_{i}^{j})\in{\mathbf{R}}_{+}^{|M|},&i=1,\dots,N,\;j=1,\dots,J,\\ &\alpha_{i}^{j}\left(\big\langle(A_{0},b_{0}),\,(G_{i}^{j},F_{i}^{j})\big\rangle+c_{0}\right)\leq 0,&i=1,\dots,N,\;j=1,\dots,J.\end{array}$
$=\begin{array}[t]{lll}\textnormal{maximize}&\sum_{i=1}^{N}\sum_{j=1}^{J}\alpha_{i}^{j}\big\langle(A_{\mathrm{obj}}^{j},b_{\mathrm{obj}}^{j}),\,(G_{i}^{j},F_{i}^{j})\big\rangle\\ \textnormal{subject to}&\sum_{i=1}^{N}\sum_{j=1}^{J}\alpha_{i}^{j}\big\|(G_{i}^{j},F_{i}^{j})-(\widehat{G}_{i},\widehat{F}_{i})\big\|\leq\varepsilon\end{array}$
with variables $(G_{i}^{j},F_{i}^{j})\in\mathcal{F}\cap\mathcal{X}$ and $(\alpha_{i}^{j})\in\Gamma$. The equality holds due to the following reasons: The case of $\alpha_{i}^{j}=0$ allows $(G_{i}^{j},F_{i}^{j})$ to be any value in ${\mathbf{S}}^{K+2}_{+}\times{\mathbf{R}}^{K+1}$, with the same objective value. Therefore, even with the constraint $\mathcal{S}(G_{i}^{j},F_{i}^{j})\in{\mathbf{R}}_{+}^{|M|}$ and $\big\langle(A_{0},b_{0}),(G_{i}^{j},F_{i}^{j})\big\rangle+c_{0}\leq 0$, the optimal value remains the same as these constraints are feasible with $(G_{i}^{j},F_{i}^{j})=(\widehat{G}_{i},\widehat{F}_{i})\in{\mathbf{S}}_{+}^{K+2}\times{\mathbf{R}}^{K+1}$.
Finally, we conclude the proof with showing the strong duality. For every $\varepsilon>0$, the minimization problem satisfies Slater’s condition if there exists a feasible point with $\lambda=0$, due to following construction of strictly feasible solution: let $\lambda=(K+2)\varepsilon$ and $(X_{i}^{j},Y_{i}^{j})=(-\varepsilon I,0)\in{\mathbf{S}}^{K+2}\times{\mathbf{R}}^{K+1}$. If $y^{j}_{\star}\in{\mathbf{R}}_{+}^{|M|}$ and $\tau^{j}_{\star}\in{\mathbf{R}}_{+}$ satisfy
$\mathcal{S}(y^{j}_{\star})+\tau^{j}_{\star}(A_{0},b_{0})-(A_{\mathrm{obj}}^{j},b_{\mathrm{obj}}^{j})\in{\mathbf{S}}_{+}^{K+2}\times\{0\},$
this leads to
$\mathcal{S}(y^{j}_{\star})-(X^{j}_{i},Y^{j}_{i})+\tau^{j}_{\star}(A_{0},b_{0})-(A_{\mathrm{obj}}^{j},b_{\mathrm{obj}}^{j})\in{\mathbf{S}}_{++}^{K+2}\times\{0\}.$
Furthermore, $\big\|(X_{i}^{j},Y_{i}^{j})\big\|=\sqrt{K+2}\varepsilon<(K+2)\varepsilon=\lambda$. Therefore, if $\varepsilon>0$, we can always construct a strictly feasible solution $y_{i}^{j}=y^{j}_{\star}$, $\tau_{i}^{j}=\tau^{j}_{\star}$, and $(X_{i}^{j},Y_{i}^{j})=(-\varepsilon I,0)$. And indeed, such feasible $y^{j}_{\star}$ and $\tau^{j}_{\star}$ exist, as they come from the solution of the dual PEP problem for the worst-case analysis. ∎
We now prove Theorem 2.
Proof of Theorem 2.
According to Lemma 3, the worst-case expectation problem is equivalent to the minimization problem in Theorem 2. Furthermore, Lemma 4 states that the minimization problem and the maximization problem in Theorem 2 are equivalent, with strong duality. ∎
We conclude this subsection with the proof of Corollary 2.1.
Proof of Corollary 2.1.
First, let $J=1$ with $(A_{\mathrm{obj}}^{1},b_{\mathrm{obj}}^{1},c_{\mathrm{obj}}^{1})=(A_{\mathrm{obj}},b_{\mathrm{obj}},0)$, which corresponds to $\mathcal{L}=\{\phi^{K}(G,F)\}=\{\big\langle(A_{\mathrm{obj}},b_{\mathrm{obj}}),(G,F)\big\rangle\}$. Using Theorem 2 automatically gives $\alpha_{i}=1/N$ for all $i=1,\dots,N$ and results in
$\sup_{\mathbf{Q}_{Z}\in B_{{\varepsilon}}}\mathop{\bf E{}}_{(G,F)\sim\mathbf{Q}_{Z}}\big(\phi^{K}(G,F))$
$=\begin{array}[t]{ll}\textnormal{minimize}&(1/N)\sum_{i=1}^{N}s_{i}\\ \textnormal{subject to}&-c_{0}\tau_{i}-\big\langle(X_{i},Y_{i}),\,(\widehat{G}_{i},\widehat{F}_{i})\big\rangle+\lambda\varepsilon\leq s_{i}\\ &-\mathcal{S}^{*}(y_{i})-(X_{i},Y_{i})+\tau_{i}(A_{0},b_{0})-(A_{\mathrm{obj}},b_{\mathrm{obj}})\in{\mathbf{S}}_{+}^{K+2}\times\{0\}\\ &\big\|(X_{i},Y_{i})\big\|_{*}\leq\lambda,\hfill i=1,\dots,N,\end{array}$
$=\begin{array}[t]{ll}\textnormal{maximize}&\sum_{i=1}^{N}\big\langle(A_{\mathrm{obj}},b_{\mathrm{obj}}),(G_{i},F_{i})\big\rangle\\ \textnormal{subject to}&\sum_{i=1}^{N}\big\|(G_{i},F_{i})-(\widehat{G}_{i},\widehat{F}_{i})\big\|\leq\varepsilon,\end{array}$
with primal variables $s_{i}\in{\mathbf{R}}$, $\tau_{i}\in{\mathbf{R}}_{+}$, $(X_{i},Y_{i})\in{\mathbf{S}}^{K+2}\times{\mathbf{R}}^{K+1}$, and $\lambda\in{\mathbf{R}}_{+}$ and dual variables $(G_{i},F_{i})\in\mathcal{F}\cap\mathcal{X}$.
Next, let $J=2$ with $\ell(G,F)=\max\{\ell^{1}(G,f),\,\ell^{2}(G,F)\}$ such that
$\left(A_{\mathrm{obj}}^{1},b_{\mathrm{obj}}^{1},c_{\mathrm{obj}}^{1}\right)=\left(\alpha^{-1}A_{\mathrm{obj}},\alpha^{-1}b_{\mathrm{obj}},-(\alpha^{-1}-1)\right),\qquad\left(A_{\mathrm{obj}}^{2},b_{\mathrm{obj}}^{2},c_{\mathrm{obj}}^{2}\right)=(0,0,1),$
gives
$\ell^{1}(G,F)=\alpha^{-1}\phi^{K}{(G,F)}-(\alpha^{-1}-1)t,\qquad\ell^{2}(G,F)=t.$
Applying Theorem 2 results in
$\begin{array}[t]{ll}\underset{\mathbf{Q}_{Z}\in B_{{\varepsilon}}}{\sup}&\mathbf{CVaR}_{\alpha}\big(\phi^{K}(G,F)\big)\end{array}$
$=\begin{array}[t]{ll}\textnormal{minimize}&(1/N)\sum_{i=1}^{N}s_{i}\\ \textnormal{subject to}&-(\alpha^{-1}-1)t-c_{0}\tau_{i}^{1}-\big\langle(X_{i},Y_{i}),(\widehat{G}_{i},\widehat{F}_{i})\big\rangle+\lambda\varepsilon\leq s_{i}\\ &t-c_{0}\tau_{i}^{2}-\big\langle(X_{i}^{2},Y_{i}^{2}),(\widehat{G}_{i},\widehat{F}_{i})\big\rangle+\lambda\varepsilon\leq s_{i}\\ &-\mathcal{S}^{*}(y_{i}^{1})-(X_{i}^{1},Y_{i}^{1})+\alpha^{-1}\tau_{i}^{1}(A_{0},b_{0})-\alpha^{-1}(A_{\mathrm{obj}},b_{\mathrm{obj}})\in{\mathbf{S}}_{+}^{K+2}\times\{0\}\\ &-\mathcal{S}^{*}(y_{i}^{2})-(X_{i}^{2},Y_{i}^{2})\in{\mathbf{S}}_{+}^{K+2}\times\{0\}\\ &\big\|(X_{i}^{1},Y_{i}^{1})\big\|_{*}\leq\lambda,\quad\big\|(X_{i}^{2},Y_{i}^{2})\big\|_{*}\leq\lambda,\qquad(i=1,\dots,N,\;j=1,\dots,J),\end{array}$
$=\begin{array}[t]{ll}\textnormal{maximize}&\alpha^{-1}\sum_{i=1}^{N}\alpha_{i}\big\langle(A_{\mathrm{obj}},b_{\mathrm{obj}}),(G_{i},F_{i})\big\rangle\\ \textnormal{subject to}&\sum_{i=1}^{N}\alpha_{i}\big\|(G_{i},F_{i})-(\widehat{G}_{i},\widehat{F}_{i})\big\|\leq\varepsilon,\end{array}$
with primal variables $s_{i}\in{\mathbf{R}}$, $t\in{\mathbf{R}}$, $\tau_{i}^{1},\tau^{2}_{i}\in{\mathbf{R}}_{+}$, $y_{i}^{1},y_{i}^{2}\in{\mathbf{R}}_{+}^{|M|}$, $(X_{i}^{1},Y_{i}^{1}),(X_{i}^{2},Y_{i}^{2})\in{\mathbf{S}}^{K+2}\times{\mathbf{R}}^{K+1}$, and $\lambda\in{\mathbf{R}}_{+}$ and dual variables $(\alpha_{i})\in\Gamma=\{(\alpha_{i})\in[0,1]^{N}\mid(1/N)\sum_{i=1}^{N}\alpha_{i}=\alpha\}$ and $(G_{i},F_{i})\in\mathcal{F}\cap\mathcal{X}$. ∎
Appendix D Numerical experiment timing results
In this section, we provide plots of all solve times in seconds for each experiment in Section 4. For the $\mathbf{CVaR}$ and expectation solve times, we report the average taken over all values of $\varepsilon$.
Table 10: Table of solve times (s) for the experiment in Section 4.1 on smooth strongly-convex quadratic minimization. Corresponds to Figure 3.
algorithm $K$ worst-case $\mathbf{CVaR}$ expectation
[…]
, late after last line=
[…]
Table 11: Table of solve times (s) for experiment in Section 4.1 on smooth non-strongly convex quadratic minimization. Corresponds to Figure 4.
algorithm $K$ worst-case $\mathbf{CVaR}$ expectation
[…]
, late after last line=
[…]
Table 12: Table of solve times (s) for experiment in Section 4.1 on smooth non-strongly convex quadratic minimization to compare with theoretical results for the MP distribution. Corresponds to Figure 5.
algorithm $K$ worst-case expectation
[…]
, late after last line=
[…]
Table 13: Table of solve times (s) for experiment in Section 4.2 on regularized logistic regression. Corresponds to Figure 6.
algorithm $K$ worst-case $\mathbf{CVaR}$ expectation
[…]
, late after last line=
[…]
Table 14: Table of solve times (s) for experiment in Section 4.3 on the Lasso problem. Corresponds to Figure 7.
algorithm $K$ worst-case $\mathbf{CVaR}$ expectation
[…]
, late after last line=
[…]